### v1

# Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory

Abstract
National Forest Inventory (NFI)s serve as the primary source of forest information, providing crucial tree species distribution data for carbon storage estimation and biodiversity assessments. However, maintaining these inventories requires labor-intensive on-site campaigns by forestry experts to identify and document tree species. Remote sensing approaches, particularly when combined with machine learning, offer opportunities to update NFIs more frequently and at larger scales. While the use of Satellite Image Time Series (SITS) has proven effective for distinguishing tree species through seasonal canopy reflectance patterns, current approaches rely primarily on Random Forest (RF) classifiers with hand-designed features and phenology-based metrics. Using deep features from an available pre-trained remote sensing foundation models offers a complementary strategy. These pre-trained models leverage unannotated global data and are meant to used for general-purpose applications and can then be efficiently fine-tuned with smaller labeled datasets for specific classification tasks. This work systematically investigates how deep features improve tree species classification accuracy in the Netherlands with few annotated data (1,462 pure species plots). We evaluate this question using three tree species datasets of varying difficulty: First the unbalanced Dutch NFI data with 13 species classes, second, unbalanced with 7 species classes, and, third, a balanced NFI dataset with 7 classes used by a RF algorithm used by recent related work. Data-wise, we extracted time-series data from Sentinel-1 (S1), Sentinel-2 (S2) and European Centre for Medium-Range Weather Forecasts Re Analysis v5 (ERA5) satellites data (January-December 2020) and Shuttle Radar Topography Mission (SRTM) data using Google Earth Engine. Our results demonstrate that fine-tuning a publicly available remote sensing time series foundation model outperforms the current state-of-the-art in NFI classification in the Netherlands by a large margin of up to 10% across all datasets. This demonstrates that classic hand-defined harmonic features are too simple for this task and highlights the potential of using deep AI features for data-limited application like NFI classification. By leveraging openly available satellite data and pre-trained models, this approach significantly improves classification accuracy compared to traditional methods and can effectively complement existing forest inventory processes.
keywords:
National Forest Inventory , deep learning , pre-trained remote sensing foundation model , time series classification
journal: International Journal of Applied Earth Observation and Geoinformation
[…]
[first]organization=Wageningen University,addressline=Droevendaalsesteeg 3 , city=Wageninegn, postcode=6708 PB, country=The Netherlands
[…]
[second]organization=OpenGeoHub Foundation,addressline=Waldeck Pyrmontlaan 14, city=Doorwerth, postcode=6865 HK, country=The Netherlands
[…]
[third]organization=Wageningen Environmental Research,addressline=P.O. Box 47, city=Wageningen, postcode=6700 AA, country=The Netherlands
1 Introduction
Forests play a significant role in mitigating climate change, adopting disaster prevention strategies, and providing ecosystem services, including sequestering carbon dioxide (CO2), providing wood materials, and serving as a source of biodiversity (Tomppo et al., 2010; FAO, 2020; Francini et al., 2024). Tree species diversity improves productivity of materials and resistance to natural disturbance (Jactel et al., 2017). Therefore, monitoring and tracking records of forests including spatial distribution of tree species are necessary for a sustainable forest management. National Forest Inventory (NFI)s are the primary source of information for various purposes such as sustainable forest management, industry investment planning, biodiversity monitoring, and Greenhouse Gas accounting (Tomppo et al., 2010; Bonannella, 2024). The Paris Agreement also requires the submission of a national carbon inventory, including carbon removals and reductions from forest lands (UNFCCC, 2015). Spatial tree species distribution information plays an important role in NFIs for various applications such as carbon storage estimation, forest management, and biodiversity assessments (Hermosilla et al., 2022; Blickensdörfer et al., 2024). Additionally, detailed tree species information is essential for national reports to the Forest Resource Assessment (FRA) of the Food and Agriculture Organization (FAO) and Forest Europe (Schelhaas et al., 2014). One of the key challenges in managing forest inventory is frequency and scale. Traditional inventory methods are based on sample-based field measurements conducted every 5 to 10 years (Tomppo et al., 2010). Current climate change and land-use changes due to economic growth cannot be captured in a timely manner using these traditional methods (Bonannella, 2024). Remote sensing helps improve and update NFIs (Hermosilla et al., 2022; Francini et al., 2024). Satellite sensors capture information about Earth every few days to several weeks, allowing us to monitor Earth’s condition more frequently and extensively, even in remote areas. However, the global coverage and complex data volumes in remote sensing make human visual interpretation impractical for large-scale analysis. Machine learning techniques have achieved significant improvements in Earth observations, including tree species classification, particularly when using multiple timestamps of images over large areas (Blickensdörfer et al., 2024; Francini et al., 2024; Hermosilla et al., 2022). The RF algorithm is one of the most popular machine learning algorithms for tree species classification due to its robustness, interpretability, and ability to handle high-dimensional data (Breiman, 2001). However, RF models require well-designed input features engineering, and this selection of appropriate features is crucial for model performance (Heaton, 2016). These choices depend on domain knowledge and target area characteristics such as climate and tree species variety, and often fail to include all necessary features (Ahlswede et al., 2023). In the context of NFIs, several state-of-the-art machine learning models utilizing RF have emerged (Hermosilla et al., 2022; Blickensdörfer et al., 2024; Francini et al., 2024)). These current methods for tree species classification rely on country-specific knowledge for input features, parameter settings, and high-quality data. However, scaling these methods to other countries at a national level presents challenges due to cost constraints and data availability limitations. Furthermore, deep learning models, which generally require large datasets to enhance performance, have remained largely unexplored for national-level tree species classification, primarily due to the limited size of available training datasets. These constraints collectively hinder the transferability of existing methods across different regions or countries. Deep learning models such as transformer architectures (Vaswani et al., 2023) have been recently introduced for forest monitoring, including tree species classification. This adoption is driven by increasing interest in multimodal and time-series data fusion in Remote Sensing (RS), enabled by the availability of big data and advancements in deep learning models (Li et al., 2022). Deep learning models can capture complex patterns in input data and create deep features that can be used for downstream tasks with classifiers and regressors including RF (Basu et al., 2015). While recent studies have achieved success in regional-scale tree species classification using high-quality labeled data, significant gaps remain in large-scale classifications (Fassnacht et al., 2016) for NFIs due to limited labeled data availability and high computational cost for training. Freely available pre-trained models, trained on large unlabeled datasets containing millions of pixels or images, have emerged as powerful tools for various downstream tasks. These models can achieve comparable or superior accuracy to traditional state-of-the-art machine learning approaches through fine-tuning without computationally expensive pre-training (Bommasani et al., 2022). Self-supervised learning, where models are trained without labels, has gained particular attention in RS applications (Wang et al., 2022). Using self-supervised learning as a model backbone with fine-tuning on limited labeled data has demonstrated significant accuracy improvements (Yu et al., 2022), particularly in time-series analysis tasks, while requiring less inductive bias (Dosovitskiy et al., 2021). The rapid growth of both labeled and unlabeled datasets for RS (Gorelick et al., 2017; Ahlswede et al., 2023) has enabled the development of various pre-trained models for tasks including tree classification (Lu et al., 2024). However, research comparing performance between fine-tuned pre-trained models and traditional approaches for tree species classification in NFIs remains limited.
This paper evaluates the effectiveness of deep features from freely available fine-tuned pre-trained models compared to a recently published random forest approach by Francini et al. (2024), who set a new state-of-the-art for Tree Species Identification in the Netherlands through a RF fitted on median reflectances and harmonic features.
The paper is structured as follows: section˜2 details the three datasets from the Dutch Forest Inventory used in this study. Section˜3 details the deep features we fine-tuned together with the harmonic features, as proposed by prior work.
2 Data
We use three datasets from the Dutch National Forest Inventory (NFI) at different complexity levels that we will use to compare the proposed fine-tuned deep time series features with the RF by Francini et al. (2024) fitted on harmonic time series features, as detailled later in section˜3.
Complex & imbalanced (COMB)
13 tree species classes, imbalanced distribution, 1,462 time series samples.
Simplified & imbalanced (SIMB)
7 tree species classes, imbalanced class distribution, 1,479 time series samples.
Simplified & balanced (SIBA)
7 species classes, a balanced class distribution, by Francini et al. (2024). In total 13,790 samples and contains additional samples augmented by photointerpretation.
Table˜1 shows a detailed list of the tree species across all three datasets.
Table 1: Species count and grouping in NFI and Francini datasets.
Dominant Species COMB Aggregated SIMB SIBA
Pinus sylvestris 513 Pinus 603 1,970
Other Pinus 89
Larix spp 56 Larix 56 1,970
Quercus robur petraea 255 Quercus 288 1,970
Other Quercus 33
Fagus spp 58 Beech 58 1,970
Populus spp 72 Populus 72 1,970
Alnus spp 30 Other Broadleaves 242 1,970
Betula spp 58
Fraxinus spp 40
Other broadleaved 102
Pseudotsuga menziesii 90 DarkConifer 160 1,970
Picea spp 66
total 1,462 1,479 13,790
2.1 Study area
This study focuses on forests in the Netherlands, which were described in the annual Greenhouse Gas accounting report (Arets et al., 2023). According to the latest National Forest Inventory 7, forests covered 363,801 ha in 2021, corresponding to 11% of the land use in the Netherlands (Schelhaas et al., 2022) and comprising approximately 36.4 million 10 $\times$ 10m pixels, as shown in fig.˜1(d).
(a) COMB (13 classes & imbalanced)
(b) SIMB (7 classes & imbalanced)
(c) SIBA (7 classes & balanced)
(d) Forests in the Netherlands
Figure 1: Overview of the study area and the three datasets used in this study. The simplified, class-balanced dataset is identical to related work by Francini et al. (2024).
2.2 National Forest Inventory
Label Data Origin Ground truth data were collected manually through field measurements at 3,062 plots for the Dutch National Forest Inventory 6 between 2012 and 2013. Each plot contained a circular area with a variable radius (5 to 20 m) to ensure inclusion of at least 20 trees (Schelhaas et al., 2022). These plots correspond to 1 to 16 pixels at a $10\times 10$m resolution of Sentinel-1 and Sentinel-2. Due to privacy considerations, the precise coordinates of the plots were obtained under a confidentiality agreement with the Dutch government, with the requirement that the data be discarded after project completion. Prior to data disposal, we utilized the plot center coordinates from the total 3,062 pixel-level data points to extract satellite data at $10\times 10$ m pixel resolution from Google Earth Engine (GEE) (Gorelick et al., 2017).
Data Quality Filtering. Following the identical protocol to Francini et al. (2024), we selected plots where a single dominant species represented more than 80% of the Basal Area (BA), which is the cross-sectional area of trees at breast height. This process removed the Castanea spp dominant species class that had no samples at this threshold.
Class Aggregation The original NFI data contains 19 dominant tree species classes. However, we noticed that six dominant species classes have less than 10 samples and we decided to remove these classes to not extenstively rely in our evaluation on underrepresented tree species, similar to Kang et al. (2017). This resulted in the dataset Complex & Imbalanced (COMB) with 1,462 samples with 13 classes for the dominant species classification task. For the Simplified & Imbalanced (SIMB), we further aggregated the classes into the 7 tree species identical to Francini et al. (2024). This resulted in 1479 data samples.
Finally, we also acquired the identical dataset used by Francini et al. (2024) which contains 13,790 data points evenly distributed across the aggregated classes (1,970 points per class). This dataset originated from the same NFI data but was augmented with additional labeled data points through visual interpretation of satellite imagery. We call Simplified & Balanced (SIBA) compared to aforementioned ones, as it contains 7 tree species classes and an even number of samples per class.
2.3 Remote Sensing Time Series
Since mono-temporal data cannot adequately capture seasonal leaf phenology for tree species classification, we utilized time-series data spanning January through December 2020. This temporal range aligns with both the Pretrained Remote Sensing Transformer (Presto) model requirements (Tseng et al., 2024) and previous Dutch tree species classification research (Francini et al., 2024).
Input Data Modalities S2, S1, ERA5, SRTM. This study employs a comprehensive multi-source remote sensing (RS) approach that integrates satellite imagery and environmental data to characterize forest vegetation across space and time. The model relies on several key data sources: Sentinel-1 SAR GRD data with VH and VV polarizations at a 6-day revisit frequency; Sentinel-2 multispectral Level 1C data with 5–10 day revisit frequency, including 10 spectral bands (B2–B8, B8A, B11, B12) and NDVI computed during preprocessing; ERA5 climate data providing monthly measurements of 2m air temperature and total precipitation; and static SRTM terrain data comprising elevation and slope.
Remote Sensing Data Preparation For S2 data, cloud masking was performed using the S2 cloud probability dataset (Pasquarella et al., 2023), excluding pixels with cloud cover probability exceeding 65% (Francini et al., 2024). Geographic coordinates were utilized both for satellite data extraction from the GEE archive and as input features for the model. Raw time-series data was downloaded as individual CSV files for each plot location across all datasets (NFI, Francini, and non-labeled data), ensuring direct correspondence between remote sensing observations and ground truth measurements. We calculated monthly median values for each band in S1 and S2 except monthly ERA5 data. After getting monthly median values for S1 and S2, Normalized Difference Vegetation Index (NDVI) was calculated from S2 bands to capture vegetation dynamics.
3 Methods
In this study, we fundamentally compare results between a classic machine learning pipeline using a RF fitted on harmonic features and seasonal reflectances by Francini et al. (2024) with pre-trained deep-learning features obtained by fine-tuning the Presto time series foundation model proposed by Tseng et al. (2024). Hence, we outline the methodological details in both complimentary approaches in the next two sections before detailing the experimental setup in section˜4.
3.1 Harmonic and Seasonal Features
We reproduced closely the classification approach of Francini et al. (2024), who recently published a new state-of-the-art classification approach for tree species mapping in the Netherlands. Their approach employ a RF Classifier (Breiman, 2001) with 500 trees that is fitted on 209 features containing Seasonal Medoid Statistics and Harmonic Features that we describe in the next two paragraphs. While Francini et al. (2024) used only features extracted from Sentinel-2 time series, we found experimentally (detailed later in section˜5.1 and table˜2) that including features from Sentinel-1 further slightly improves the results. Since we will also use Sentinel-1 to extract deep features (section˜3.2), we also include Sentinel-1 features to ensure a fair comparison.
76 Seasonal Medoid Statistics
In total, 12 bands from S1 and S2 are selected and additional 7 indices are calculated from S2 including NDVI, Normalized Burn Ratio (NBR), Enhanced Vegetation Index (EVI), and Tasseled Cap transformations: Brightness (TCB), Wetness (TCW), Greenness (TCG), Angle (TCA). This results in 19 bands, where for each band the median band value in each of the 4 seasons, winter (January, February, December), spring (March-May), summer (June-August), and autumn (September-November), the medoids values are extracted. These features describe the data through season-wise statistics, but do not capture the dynamic change in seasonality itself.
133 Harmonic Features
To capture seasonal changes, sine and cosine harmonics
$P_{t}=\beta_{0}+\beta_{1}t+\beta_{2}\cos(2\pi\omega t)+\beta_{3}\sin(2\pi\omega t)$ (1)
are fitted to the underlying monthly signal of each band according. The seven harmonic parameters $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $A$, $\phi$ for each of the 19 band are used as classification features alongside the residual RMSE: $\beta_{0}$ is the constant, $\beta_{1}$ is the time coefficient, and $\beta_{2}$ and $\beta_{3}$ are the frequency sine and cosine coefficients, respectively. To fit these four coefficients and to select $P_{t}$, the pixel $p$ harmonic values at time $t$, we used a least squares regression to fit Eq. 1. $A=\sqrt{\beta_{2}^{2}+\beta_{3}^{2}}$ is the amplitude of the harmonic curve on the y-axis, $\phi=\arctan(\beta_{3}/\beta_{2})$ is the phase of the curve on the x-axis to the origin and $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(P_{t}-X_{t})^{2}}$ is root mean square error between $P_{t}$ and the actual pixel values $X_{t}$. The frequency $\omega$ = 1 (in years) and indicates one cycle per unit of time, initial guess values before fitting four coefficients were set as $\beta_{0}=0.1$, $\beta_{1}=0.1$, $\beta_{2}=0.4$, $\beta_{3}=0.4$.
This approach was proposed by Francini et al. (2024) based on Sentinel-2 features. We added of Sentinel-1 features, as we found them to be slightly beneficial in accuracy (see section˜5.1).
3.2 Deep Pre-trained Time Series Features
Overview of Presto deep features. In contrast to harmonic features that fit mathematical curves to individual bands, we evaluate the effectiveness of deep features obtained by the pre-trained time series foundation model, which is Lightweight, Pre-trained Transformers for Remote Sensing Timeseries (Presto) developed and trained by Tseng et al. (2024) for general-purpose remote sensing time series classification and regression. This model is based on transformer architecture and was originally pretrained on 21.5 million pixel time-series with 12-month contiguous intervals and static topographic and coordinates data. Each month’s composite satellite data included S1, S2 and its NDVI, ERA5, Dynamic World (DW), extracted between 2020-01-01 and 2021-12-31. The static SRTM and coordinates data are added to the time-series data as input of the Presto encoder. The model was originally pretrained through Masked Auto Encoding (MAE) (He et al., 2022) reconstruct masked sequences across different satellite modalities. The advantages of this model include the ability to handle freely available multi-source and multi-temporal data even when some sources (e.g., ERA5) or temporal data (e.g., S2 values in November) are missing, and computational efficiency when processing large areas. We obtained 128 dimensional deep features from the encoder output.
For this study, we used the pre-trained weights to initialize the Presto Encoder.
Multi Layer Perceptron (MLP) classifier was attached for fine-tuning of Presto encoder to the downstream task of tree species classification, which is described in section˜4. We used the fine-tuned encoder to extract deep features from the test set for the comparison between Harmonic and Seasonal Features and Presto deep features.
As shown in fig.˜2, we directly input 12-monthly time series ($T=12$ timesteps) for (i) 2 Sentinel-1 bands (S1), (ii) 10 Sentinel-2 bands (S2) and (iii) NDVI (NDVI), (iv) 2 ERA5 features (ERA5), (v) Dynamic world tensor with all tree classes value (equal to value of 1. DW). To every pixel-time series we appended two static-in-time products: (i) elevation and slope from the SRTM digital elevation model (Farr et al., 2007) (TG) and (ii) location coordinates (Loc) of each pixel as input tokens to fine-tuned Presto encoder. Location coordinates (Loc) is 3D static in time Cartesian coordinates computed from the latitude and longitude of the pixel’s geographical location.
One training sample $\mathbf{x}$, comprising a pixel-timeseries $\mathbf{t}\in\mathbb{R}^{T\times 15}$, land cover classes $\mathbf{v}\in\mathbb{V}^{T\times 1}$, and static variables $\mathbf{s}\in\mathbb{R}^{1\times 5}$, is summarized as follows:
$\mathbf{x}=\left[\{t^{\text{S1}}_{i};t^{\text{S2}}_{i};t^{\text{ERA5}}_{i};t^{\text{NDVI}}_{i};t^{\text{DW}}_{i}\,\middle|\,i=1,\ldots,12\,\};\,s^{\text{TG}};s^{\text{Loc}}\right]$ (2)
The pixel-timeseries $\mathbf{x}$ were transformed into a sequence of tokens, each represented by an embedding $\mathbf{e}$, suitable for processing by the Presto transformer. For each timestep $0\leq i<T$, input variables were split into channel groups $\mathcal{C}$ based on their sensor or data source as follows;
  • 1.
    $\mathbf{e}_{\text{S1}}$: S1 VV and VH bands.
  • 2.
    $\mathbf{e}_{\text{S2-RGB}}$: S2 RGB(B2, B3, B4) bands.
  • 3.
    $\mathbf{e}_{\text{S2-RE}}$: S2 Red Edge (B5, B6, B7) bands.
  • 4.
    $\mathbf{e}_{\text{S2-NIR10}}$: S2 NIR (B8) 10m band.
  • 5.
    $\mathbf{e}_{\text{S2-NIR20}}$: S2 NIR (B8A) 20m band.
  • 6.
    $\mathbf{e}_{\text{S2-SWIR}}$: S2 SWIR(B11, B12) bands.
  • 7.
    $\mathbf{e}_{\text{NDVI}}$: Computed from S2 B4 and B8.
  • 8.
    $\mathbf{e}_{\text{ERA5}}$: ERA5 Precipitation and 2m temperature.
  • 9.
    $\mathbf{e}_{\text{TG}}$: Elevation and slope from SRTM DEM.
  • 10.
    $\mathbf{e}_{\text{Loc}}$: Cartesian coordinates from latitude/longitude.
Each real-valued group was projected into a common latent space of dimension $d_{e}$ using group-specific learned linear projections $h_{C}$, such that $e^{\text{S1}}_{i}=h_{\text{S1}}(t^{\text{S1}}_{i})$. Categorical variables like Dynamic World classes were embedded via indexing into a learnable embedding matrix. To convey the (i) location, (ii) timestamp, and (iii) channel group of each token, specific encodings were added to the embeddings $\mathbf{e}$. The final embedding of dimension $d_{e}$ is a concatenation of:
  • 1.
    Positional Encoding: Sinusoidal encoding as in Vaswani et al. (2023).
  • 2.
    Month Encoding: To represent seasonal similarity, each month (0 to $11$) is encoded as:
    $p_{\text{month},2i}$ $=\sin\left(\frac{2\pi\cdot\text{month}}{12}\right)$
    $p_{\text{month},2i+1}$ $=\cos\left(\frac{2\pi\cdot\text{month}}{12}\right)$
For static variables, positional and month encodings were set to zero.
  • 1.
    Channel Group Encoding: A learnable encoding was added for each channel group $c\in\mathcal{C}=\{\text{S1},\text{S2 RGB},\ldots,\text{ERA5},\text{TG},\text{Loc}\}$, to capture source-specific information.
The final transformer input matrix $\mathbf{E}\in\mathbb{R}^{(T\cdot|\mathcal{C}_{\text{dynamic}}|+|\mathcal{C}_{\text{static}}|)\times d_{e}}$ is composed of:
  • 1.
    Dynamic Variables: For each timestep $i<T$ and channel group $c\in\mathcal{C}$:
    $\mathbf{e}^{c}_{i}=h_{c}(t^{c}_{i})+[p^{c}_{\text{channel}};p_{\text{sin}}(i);p_{\text{month}}(i)]$
  • 2.
    Topographical Data:
    $\mathbf{e}_{\text{TG}}=h_{\text{TG}}(s_{\text{TG}})+[p^{\text{TG}}_{\text{channel}};0;0]$
  • 3.
    Coordinates:
    $\mathbf{e}_{\text{Loc}}=h_{\text{Loc}}(s_{\text{Loc}})$
Figure 2: Presto encoder feature extraction for downstream tasks. Figure modified from (Tseng et al., 2024).
Data Normalization (for deep features). The data normalization follows the original Presto setup (Tseng et al., 2024), with adjustments based on the observed value ranges for each source. Sentinel-1 VH and VV bands, originally ranging from –31 to 17 dB, are normalized by shifting +25 and dividing by 25, resulting in a range of –0.24 to 1.68. Sentinel-2 spectral bands, with raw values from 10 to 15,769, are scaled by dividing by 10,000, yielding normalized values between 0.00 and 1.58. ERA5 temperature data, originally 278 to 295 K (equivalent to 6 to 23°C), are shifted by –272.15 and divided by 35, producing a range of 0.17 to 0.66. ERA5 precipitation, ranging from 0.008 to 0.208 m, is normalized by dividing by 0.03, resulting in values from 0.27 to 6.93. SRTM elevation data, spanning –27 to 331 m, are divided by 2,000 to give a range of –0.01 to 0.17, while slope values from 0.0 to 39.3° are scaled by dividing by 50, yielding normalized values between 0.00 and 0.79.
4 Experimental Setup
This study compares classic machine learning with hand-crafted feature engineering against deep learning approaches for tree species classification. Our experimental design follows a systematic progression to evaluate different methodological approaches and their combinations.
4.1 Comparative evaluation framework
Our experimental design implements four distinct comparison scenarios to systematically evaluate feature representations and classifier architectures:
  1. 1.
    Sensor contribution analysis: S1 vs S2 vs S1+S2 using harmonic features with RF
  2. 2.
    Feature representation comparison: Harmonic features vs deep features using RF
  3. 3.
    Classifier architecture evaluation: RF vs MLP using deep features
This systematic approach allows us to isolate the contributions of different methodological components while building toward the optimal configuration for operational NFI applications, extending the work of Francini et al. (2024) with comprehensive deep learning comparisons.
4.2 Model fine-tuning
The fine-tuning approach leverages the pre-trained Presto encoder’s learned representations while adapting the model to the specific characteristics of Dutch forest species. This transfer learning strategy balances computational efficiency with task-specific optimization.
For fine-tuning the pre-trained models, the encoder block from the pre-trained model was reused and trained with newly attached Multi Layer Perceptron (MLP) classifier for the downstream task. This fine-tuning model is trained on the 1,479 and 13,790 samples for 7 aggregated groups classification task, and the 1,462 samples for the 13 dominant species classification task. Pre-trained encoder was set as trainable in this process.
Based on the current state-of-the-art classifier of MLP on tree species classification (Mouret et al., 2024), this classifier architecture consists of a 3-layer MLP with no dropout (rate: 0.0). The layer configuration comprises 1024 nodes in the first layer, 512 nodes in the second layer, and 256 nodes for the last layer, with batch normalization and ReLU activation applied throughout. The training parameters are configured as follows: a learning rate of 0.0001 is used with a weight decay of 0.00746. The model is trained for a maximum of 100 epochs using a batch size of 64. The training process employs cross-entropy loss for optimization. The best fine-tuned models and its accuracy were selected based on best validation loss.
4.3 Data splitting strategy
Train-test split ratio is a critical factor in model performance on downstream tasks. If dataset size is large, 80/20 or higher ratios are commonly used (Rácz et al., 2021; Joseph, 2022). Based on our imbalanced dataset and previous research (Blickensdörfer et al., 2024), a 70/30 (approximately 2:1 ratio) train-test split was applied for fine-tuning and RF classification to increase model performance and reduce standard deviation on tree species classification over large areas for NFI. During fine-tuning and subsequent RF classification for features extracted from the fine-tuned model, the same training and test data were used. This practice ensures that extracted features are not trained on test data, which would cause data leakage. By adopting a class-per-split strategy, each class is equally split into train and test data at this ratio.
In remote sensing machine learning, ensuring that the train-test split strategy does not introduce spatial autocorrelation is crucial for obtaining unbiased model performance estimates Karasiak et al. (2022). Our NFI data accommodates a sampling strategy which selects each plot with a density of 1 point per 100ha (Schelhaas et al., 2014). This means that each random point is designated per square kilometer. Additionally, the Francini data removed adjacent pixels with a minimum distance of 15m (Hermosilla et al., 2022; Francini et al., 2024). These sampling strategies ensure that the train-test split does not introduce strong spatial autocorrelation in the data; therefore, no additional train-test split strategy was applied in this study.
4.4 Performance evaluation
Model performance was evaluated using a comprehensive set of established accuracy metrics: per-class precision and recall, F1 scores, confusion matrices, and overall accuracy. While overall accuracy provides a general assessment of classification performance, it exhibits reduced sensitivity to class imbalance issues that are prevalent in ecological datasets. Therefore, the F1 score, which represents the harmonic mean of precision and recall, was employed as a supplementary metric to provide a more balanced evaluation of model performance when addressing the inherent class imbalances in tree species distribution data. Confusion matrices were used to visualize the model’s performance at the class level. These metrics are widely used in tree species classification studies (Goutte and Gaussier, ; Hermosilla et al., 2022; Francini et al., 2024; Blickensdörfer et al., 2024) and provide comprehensive assessment of classification performance across all species classes.
4.5 Statistical validation
To evaluate variability of model performance, each model was evaluated five times with different random seeds, which affect to train-test split and model initialization. The mean and standard deviation of the accuracy metrics were calculated for each model. This approach provides a more robust evaluation of model performance.
5 Results
5.1 Influence of multi-source input features
Following the approach of Francini et al. (2024), we first evaluated the contribution of Sentinel-1 radar data to optical Sentinel-2 classification. Using RF with harmonic features, the addition of Sentinel-1 provided modest improvements: 0.4% for the 13-class COMB dataset and 1.4% for the 7-class SIMB dataset. These results justify the inclusion of Sentinel-1 features in our subsequent comparisons.
We first evaluated whether combining optical and radar satellite data improves tree species classification using hand-crafted features in a RF classifier. Feature sets were derived from seasonal medoid composites and harmonic fits, using either Sentinel-2 (S2) alone or in combination with Sentinel-1 (S1). The goal was to assess whether radar backscatter, which captures structural and moisture-related canopy properties, offers complementary information to optical reflectance data. As shown in Table 2 the inclusion of S1 provided modest but consistent improvements across both datasets. For the 13-class dataset (COMB), accuracy increased from 66.34% to 66.71% when adding S1 to S2-derived features. In the 7-class dataset (SIMB), accuracy improved from 72.87% to 74.22%. These results suggest that the radar signal contributes marginally to distinguishing between species, especially under a more simplified class grouping. While the absolute gains remain small, incorporating S1 ensures a more robust feature representation, particularly in settings where optical signals may be limited by cloud cover or seasonal variability. The modest accuracy gains observed here are also useful to contextualize the more substantial differences in performance reported in the next sections, which stem from the feature representations rather than the data sources themselves.
Table 2: Classification accuracy comparison across different feature combinations and satellite data sources in NFI data (in %). Bold values indicate best performance for each label type.
Label Features Sentinel-1 Sentinel-2 S1-S2 Combined
Band Acc. Band Acc. Band Acc.
COMB (13 classes) Seasonal (S) S1 51.03 S2 64.03 S1-S2 64.52
Harmonic (H) S1 55.04 S2 63.79 S1-S2 64.64
All (S+H) S1 57.72 S2 66.34 S1-S2 66.71
SIMB (7 classes) Seasonal (S) S1 55.60 S2 70.18 S1-S2 71.93
Harmonic (H) S1 60.19 S2 71.39 S1-S2 71.52
All (S+H) S1 62.75 S2 72.87 S1-S2 74.22
5.2 Deep vs Harmonic features
Table 3 demonstrates the superior performance of deep features over hand-crafted features when both are evaluated using RF classification. Deep features consistently outperform traditional harmonic and seasonal features across all three datasets, with F1 score improvements of 6-8 percentage points for the imbalanced datasets (COMB and SIMB) and over 10 percentage points for the balanced dataset (SIBA).
Table 3: Classification performance comparison (in %) across feature types (deep vs. harmonic+medoid) and classifiers (MLP vs. RF). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance among MLP and RF per dataset type.
Classification Results
Dataset type Data size MLP Classifier
(Deep features)
RF Classifier
(Deep features)
RF Classifier
(Harm.+med. Francini et al. (2024))
Overall Accuracy SIMB (7 classes) 1479 77.13 ± 1.17 74.62 ± 1.93 70.58 ± 1.78
COMB (13 classes) 1462 67.99 ± 1.33 66.77 ± 1.51 62.71 ± 1.51
SIBA (7 classes) 13,790 98.21 ± 0.20 95.28 ± 0.60 84.31 ± 0.99
F1 Score SIMB (7 classes) 1479 68.24 ± 2.85 60.54 ± 3.08 51.57 ± 3.36
COMB (13 classes) 1462 52.45 ± 2.25 46.84 ± 2.65 38.54 ± 2.14
SIBA (7 classes) 13,790 98.20 ± 0.20 95.27 ± 0.60 84.26 ± 0.99
For direct comparison using identical S1+S2 sensor inputs (detailed in Appendix Table A.1), deep features show F1 score improvements of 12.3 percentage points (SIMB), 9.7 percentage points (COMB), and 10.7 percentage points (SIBA), confirming consistent performance advantages across different input configurations.
We compared classification performance between traditional hand-crafted features and deep features extracted from a pre-trained transformer model. In both cases, a RF classifier was used to isolate the effect of the feature representation. As shown in Table 3, deep features extracted from the pre-trained Presto model consistently outperformed the harmonic and medoid predictors across all dataset.
On the 7-class dataset (SIMB), models trained with deep features achieved an overall accuracy of 74.62% (±1.93), compared to 70.58% (±1.78) for the harmonic+medoid approach – an improvement of approximately 4 percentage points. The difference was even more pronounced when examining F1 scores, where deep features yielded 60.54% (±3.08) versus 51.57% (±3.36) for traditional features, representing an improvement of nearly 9 percentage points. In the more complex 13-classes case (COMB), deep features kept the advantage over traditional hand-crafted features, achieving an overall accuracy of 66.77% (±1.51) compared to 62.71% (±1.51). Similarly, the F1 score showed a substantial improvement from 38.54% (±2.14) with traditional features to 46.84% (±2.65) with deep features – a gain of more than 8 percentage points. This indicates that the deep learning approach handles the increased complexity of multi-class classification more effectively. The highest performance gains were observed in the Francini dataset (SIBA), where deep features achieved a remarkable overall accuracy of 95.28% (±0.60), representing an improvement of nearly 11 percentage points over the traditional features’ accuracy of 84.31% (±0.99). The F1 scores demonstrated comparable improvements, with deep features achieving 95.27% (±0.60) compared to 84.26% (±0.99) for traditional features.
These results suggest that deep features extracted from the pre-trained transformer model are more effective at capturing species-specific spectral and temporal signatures than traditional hand-crafted features. The improvement is most pronounced in the SIBA dataset – where class separation is clearer due to the balanced classes and a higher sample size per class – but performance gains are evident across all dataset configurations examined in this study.
To evaluate the influence of classifier architecture, we compared the Multi Layer Perceptron (MLP) classifier for tree species classification Mouret et al. (2024) with a RF, both trained on the same deep features extracted from the pre-trained Presto model. This isolates the effect of the classifier while keeping the input representation fixed.
Across all datasets, the MLP classifier consistently outperformed the RF classifier in both overall accuracy and F1 score (Table 3). In the the 7-class dataset (SIMB), the MLP achieved an overall accuracy of 77.13% (±1.17) compared to 74.62% (±1.93) for the RF with a difference of approximately 2.5 percentage points. The performance gap was even more pronounced in terms of F1 score, with the MLP achieving 68.24% (±2.85) versus 60.54% (±3.08) for RF, representing a substantial improvement of nearly 8 percentage points. For the more challenging 13-class dataset (COMB), the MLP classifier maintained its advantage with an overall accuracy of 67.99% (±1.33) compared to 66.77% (±1.51) for the RF classifier. Similarly, the F1 score showed the MLP’s superior performance with 52.45% (±2.25) versus 46.84% (±2.65) for RF. The largest performance gain was observed in the balanced dataset (SIBA), where MLP reached 98.21% accuracy and 98.20% F1, compared to 95.28% and 95.27% with RF.
Notably, the MLP classifier not only achieved higher mean performance values but also demonstrated more consistent results across repeated runs, as evidenced by the lower standard deviations in most cases. These results indicate that while both classifiers can effectively leverage deep features, the MLP classifier with pre-trained Presto offers superior performance for forest type classification tasks across datasets of varying complexity and size. The improved class separation suggests that MLPs better exploit species-specific spectral and temporal signatures embedded in the deep features—patterns that may be less accessible to tree-based models like RF, especially in more balanced or data-rich scenarios.
The S1+S2 only comparison (detailed in Appendix Table A.2) confirms these findings with F1 score improvements of 6.7 percentage points for SIMB (70.57% vs 63.83%), 6.7 percentage points for COMB (54.91% vs 48.24%), and 3.1 percentage points for SIBA (98.07% vs 94.94%), demonstrating consistent MLP advantages across different input configurations.
5.3 Class level performances
To better understand class-level behavior, we visualized the confusion matrices of the best-performing model – an MLP fine-tuned on deep features – for all three datasets (COMB, SIMB, SIBA). These matrices are shown in Figure 3, with each cell reporting the number of predicted samples and the corresponding precision (i.e., percentage of correct predictions among all predicted samples for that class). In the 7-class NFI dataset (SIMB), the model performs strongly across most classes. Pinus shows high precision with 1,668 correctly predicted samples and a precision of 84.7%. Dark Conifer and Populus also achieve high precision with 1,448 samples at 73.5% and 1,833 samples at 93.0%, respectively. Broadleaved classes exhibit more overlap: Quercus (oak) has 1,478 samples with 75.0% precision, while Other Broadleaves and Fagus reach 71.6% and 91.6%, respectively. Confusion is mostly observed between Quercus, Populus, and Other Broadleaves, reflecting the spectral similarity within this group.
In the more complex 13-class dataset (COMB), misclassifications are more pronounced, but key species maintain strong separation. Pinus sylvestris achieves 77.4% precision, Pseudotsuga menziesii reaches 83.3%, and Picea spp. also stands at 83.3%. Populus spp. improves to 73.9%, and Betula spp. achieves 100% precision, though on a smaller sample size. In contrast, Quercus robur/petraea reaches only 53.9%, and Other Broadleaves suffers from overlapping predictions with Quercus and Populus. Despite the increased class count, precision remains relatively high for most coniferous classes and for broadleaves with distinct phenology.
In the balanced dataset (SIBA), the model achieves near-perfect class separation. All classes have 1,970 samples, and precision ranges between 84.0% (Larix) and 93.0% (Populus). Other classes include Fagus (91.6%), Pines (84.7%), Oak (75.0%), and Other Broadleaves (71.6%). The confusion is minimal and symmetrically distributed, confirming that the model benefits substantially from balanced data and larger sample size.
Figure 3: Confusion matrices for the best performing models with deep features by MLP classifier. Simplified & Imbalanced (SIMB)(left), Complex & Imbalanced (COMB)(center), Simplified & Balanced (SIBA) (right)
6 Discussion
6.1 Multi-modal sensor integration
Our findings on the modest but consistent benefits of combining Sentinel-1 radar with Sentinel-2 optical data align well with recent European tree species mapping studies. Blickensdörfer et al. (2024) demonstrated that Sentinel-1 SAR data effectively fills potential data gaps in cloud-affected regions, particularly important for national-scale applications where clear-sky observation density varies substantially. Similarly, Liu et al. (2023) confirmed that Sentinel-1 provides supplementary structural information to Sentinel-2 optical data, contributing to increased reliability of large-scale tree species mapping. Our results showing 0.4-1.4% accuracy improvements when adding Sentinel-1 to Sentinel-2 features corroborate these findings and provide confidence in the robustness of multi-modal approaches for operational NFI applications.
6.2 Deep vs hand-crafted features: an emerging paradigm
The transition from hand-crafted to learned feature representations marks a significant shift in remote sensing methodology. While deep learning has proven transformative in crop type mapping and land cover classification (Jakubik et al., 2023; Szwarcman et al., 2025; Wang et al., 2023), its application to tree species classification at NFI scale represents a relatively recent development. Our results demonstrate consistent improvements in F1 scores of 6-12 percentage points across datasets of varying class balance and complexity, suggesting that transformer-derived features capture phenological and structural variability more effectively than engineered metrics. This performance gain supports the ongoing shift in remote sensing from task-specific feature engineering toward transferable, learned representations that can adapt to diverse ecological contexts without domain-specific tuning.
6.3 Neural network architectures for learned features
The systematic superiority of MLP classifiers over RF when both were trained on deep features reflects broader patterns observed in high-dimensional, learned feature spaces. Neural architectures tend to better exploit the latent structure present in deep representations, where complex non-linear relationships between features may not be effectively captured by tree-based models. This finding aligns with recent advances in representation learning, where the pairing of learned features with compatible downstream architectures has proven critical for optimal performance (Mouret et al., 2024).
6.4 Computational efficiency and operational scalability
From an operational perspective, this approach offers significant advantages for national-scale implementations. The computational efficiency is evidenced by Presto’s lightweight architecture, which enables fine-tuning on a single GPU or CPU (Tseng et al., 2024), making it accessible to organizations with limited computational resources. The pre-training process itself demonstrates remarkable efficiency, processing 21.5 million pixel time-series in only 2 hours 12 minutes per epoch (43 hours 15 minutes for 20 epochs), indicating that similar approaches could be readily applied to other countries’ NFI systems. For The Netherlands specifically, with its 36.4 million 10×10m pixels, large-scale pixel-label mapping appears computationally feasible using this methodology.
6.5 Data quality and class balance considerations
Our comparative analysis across datasets of varying quality and balance reveals critical insights for operational NFI implementations. The stark performance differences between the NFI datasets (COMB and SIMB) and the Francini dataset (SIBA) underscore the fundamental importance of training data characteristics for deep learning approaches.
The Francini dataset’s superior performance (98.2% F1 score) compared to the imbalanced NFI datasets (52.5-68.2% F1 scores) primarily reflects the impact of class balance rather than dataset size alone. While the Francini dataset benefits from both larger sample size (13,790 vs 1,462-1,479) and better class balance, the magnitude of improvement suggests that addressing class imbalance represents a critical optimization opportunity for operational systems.
This finding has important implications for NFI data collection strategies. Resources may be more effectively allocated to targeted collection for underrepresented species classes. The approach demonstrated by Francini et al. (2024), which supplements ground-truth NFI data with additional samples through visual interpretation of high-resolution satellite imagery, offers a practical pathway for addressing these imbalances. Tools such as Collect Earth Online (Saah et al., 2019) provide accessible platforms for additional data collection.
6.6 Limitations and future directions
Several limitations merit consideration for operational deployment. First, deep features were extracted from a model pre-trained on a limited temporal and spatial range (2020-2021 global data), and further domain adaptation might improve performance for underrepresented classes to country’s forests. Second, while MLPs demonstrated superior performance in our controlled experiments, their generalization in operational settings with smaller or more imbalanced training sets requires further investigation.
Future research directions include: (i) adaptation to new regions or tree species not present in the training data, testing the transferability of learned representations across different forest ecosystems; (ii) integration of additional modalities such as hyperspectral or LiDAR data where available, potentially capturing structural metrics that complement spectral-temporal patterns; and (iii) development of computational efficient uncertainty quantification methods to assess classification confidence for operational decision-making.
7 Summary and conclusions
This study examined the efficacy of deep learning approaches for tree species classification in National Forest Inventories, yielding several significant findings. First, deep features extracted from pre-trained Presto models consistently outperformed traditional methods employing harmonic and medoid predictors across varying dataset sizes and classification complexities. This performance advantage was particularly evident with larger and more balanced training datasets, as demonstrated by the results from the Francini dataset, indicating superior scalability of deep learning approaches with increased high-quality data availability. Second, the MLP classifier demonstrated superior performance compared to RF in our experimental setting, corroborating previous findings (Mouret et al., 2024). Collectively, these results demonstrate the potential for implementing computationally efficient, large-scale tree species mapping within NFI systems.
This research significantly advances our understanding of deep learning applications in forestry by having demonstrated the effectiveness of fine-tuning pre-trained deep learning models for tree species classification, establishing a computationally efficient approach for improving NFI accuracy.
This insight lays the groundwork for future developments in automated forest inventory systems and broader applications in environmental monitoring.
CRediT authorship contribution statement
Takayuki Ishikawa: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis, Conceptualization. Carmelo Bonannella: Writing – review & editing, Supervision, Conceptualization. Bas J. W. Lerink Writing – review & editing, Data curation. Marc Rußwurm Writing – review & editing, Supervision, Conceptualization.
Declaration of competing interest
C.B. reports employment with OpenGeoHub Foundation. Other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
This research is based on data from the Dutch National Forest Inventory (NFI), commissioned by the Ministry of Agriculture, Nature and Food Quality of the Netherlands. We thank the Ministry and the agencies responsible for collecting and providing these nationally representative forest data, which underpin our analysis. C.B. (Carmelo Bonannella) was supported by the Open-Earth-Monitor Cyberinfrastructure project, funded by the European Union’s Horizon Europe research and innovation programme [grant agreement No.101059548].
Data availability
The source code is available on:
Appendix A Appendix
A.1 Additional results
A.1.1 Results of models using only S1 and S2 data for fine-tuning
Our primary models incorporated all available satellite data as input for both the second stage of pre-training and fine-tuning to extract deep features. In this section, we present results from models using only S1 and S2 data as input for fine-tuning and feature extraction, excluding SRTM and ERA5 data.
Deep features model performance
Table A.1 demonstrates that deep features extracted by pre-trained Presto from only S1 and S2 data outperformed harmonic and medoid features in the RF classifier. These results were comparable to models that utilized all satellite data as input.
Table A.1: Classification performance comparison between deep features and harmonic+medoid features (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between deep features extracted from Presto and harmonic+medoid features per dataset type.
RF
Dataset type Data size Deep feat.
((Tseng et al., 2024))
Harm.+med.
((Francini et al., 2024))
Overall Accuracy NFI data (7 classes) 1480 75.47 ± 1.15 70.58 ± 1.78
NFI data (13 classes) 1462 67.77 ± 2.38 62.71 ± 1.51
Francini data (7 classes) 13,790 94.96 ± 0.58 84.31 ± 0.99
F1 Score NFI data (7 classes) 1480 63.83 ± 1.21 51.57 ± 3.36
NFI data (13 classes) 1462 48.24 ± 4.31 38.54 ± 2.14
Francini data (7 classes) 13,790 94.94 ± 0.58 84.26 ± 0.99
Table A.2 shows that the MLP classifier consistently achieved higher accuracy than the RF classifier, even when using only S1 and S2 data as input for the pre-trained Presto model.
Table A.2: Classification performance comparison between MLP and RF classifiers with deep features (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between MLP and RF classifier per dataset type.
Results
Dataset type Data MLP RF
size classifier classifier
Overall Accuracy NFI data (7 classes) 1480 78.79 ± 1.01 75.47 ± 1.15
NFI data (13 classes) 1462 69.39 ± 1.65 67.77 ± 2.38
Francini data (7 classes) 13,790 98.08 ± 0.09 94.96 ± 0.58
F1 Score NFI data (7 classes) 1480 70.57 ± 2.20 63.83 ± 1.21
NFI data (13 classes) 1462 54.91 ± 4.29 48.24 ± 4.31
Francini data (7 classes) 13,790 98.07 ± 0.09 94.94 ± 0.58
Comparison between all bands and only S1 and S2 bands
Table A.3 compares models using all satellite data against those using only S1 and S2 data with MLP classifier. Interestingly, models with only S1 and S2 data achieved higher accuracy with NFI data, while no significant difference was observed with Francini data.
Table A.3: Classification performance comparison between using all available bands versus using only S1 and S2 bands (in %). Each model was run five times with mean and standard deviation (±) reported. Bold values indicate best performance between the two band configurations for each dataset type.
MLP classifier
Dataset type Data size All bands S1S2 only
Overall Accuracy NFI data (7 classes) 1480 77.13 ± 1.17 78.79 ± 1.01
NFI data (13 classes) 1462 67.99 ± 1.33 69.39 ± 1.65
Francini data (7 classes) 13,790 98.21 ± 0.20 98.08 ± 0.09
F1 Score NFI data (7 classes) 1480 68.24 ± 2.85 70.57 ± 2.20
NFI data (13 classes) 1462 52.45 ± 2.25 54.91 ± 4.29
Francini data (7 classes) 13,790 98.20 ± 0.20 98.07 ± 0.09

### v2

@@
-Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory
+Assessing the Effectiveness of Deep Embeddings for Tree Species Classification in the Dutch Forest Inventory
 Abstract
-National Forest Inventory (NFI)s serve as the primary source of forest information, providing crucial tree species distribution data for carbon storage estimation and biodiversity assessments.
-However, maintaining these inventories requires labor-intensive on-site campaigns by forestry experts to identify and document tree species.
-Remote sensing approaches, particularly when combined with machine learning, offer opportunities to update NFIs more frequently and at larger scales.
-While the use of Satellite Image Time Series (SITS) has proven effective for distinguishing tree species through seasonal canopy reflectance patterns, current approaches rely primarily on Random Forest (RF) classifiers with hand-designed features and phenology-based metrics.
-Using deep features from an available pre-trained remote sensing foundation models offers a complementary strategy.
-These pre-trained models leverage unannotated global data and are meant to used for general-purpose applications and can then be efficiently fine-tuned with smaller labeled datasets for specific classification tasks.
-This work systematically investigates how deep features improve tree species classification accuracy in the Netherlands with few annotated data (1,462 pure species plots).
-We evaluate this question using three tree species datasets of varying difficulty: First the unbalanced Dutch NFI data with 13 species classes, second, unbalanced with 7 species classes, and, third, a balanced NFI dataset with 7 classes used by a RF algorithm used by recent related work.
-Data-wise, we extracted time-series data from Sentinel-1 (S1), Sentinel-2 (S2) and European Centre for Medium-Range Weather Forecasts Re Analysis v5 (ERA5) satellites data (January-December 2020) and Shuttle Radar Topography Mission (SRTM) data using Google Earth Engine.
-Our results demonstrate that fine-tuning a publicly available remote sensing time series foundation model outperforms the current state-of-the-art in NFI classification in the Netherlands by a large margin of up to 10% across all datasets.
-This demonstrates that classic hand-defined harmonic features are too simple for this task and highlights the potential of using deep AI features for data-limited application like NFI classification.
-By leveraging openly available satellite data and pre-trained models, this approach significantly improves classification accuracy compared to traditional methods and can effectively complement existing forest inventory processes.
+National Forest Inventory (NFI) serves as the primary source of forest information, however, maintaining these inventories requires labor-intensive on-site campaigns by forestry experts to identify and document tree species.
+Embeddings from deep pre-trained remote sensing models offer new opportunities to update NFIs more frequently and at larger scales.
+While training new deep learning models on few data points remains challenging, we show that using pre-computed embeddings can directly outperform existing hand-designed features from Satellite Image Time Series (SITS) that have proven effective for distinguishing tree species through seasonal canopy reflectance patternsin combination with Random Forest (RF) classifiers.
+This work systematically investigates how deep embeddings improve tree species classification accuracy in the Netherlands with few annotated data (1,462 pure species plots).
+We evaluate this question on three embedding models: Presto, Alpha Earth, and Tessera, using three tree species datasets of varying difficulty.
+Data-wise, we compare the available embeddings from Alpha Earth and Tessera with dynamically calculated embeddings from a pre-trained Presto model, for which we extracted time series from Sentinel-1 (S-1), Sentinel-2 (S-2), and weather data, along with elevation data downloaded from Google Earth Engine.
+Our results demonstrate that fine-tuning a publicly available remote sensing time series pre-trained model outperforms the current state-of-the-art in NFI classification in the Netherlands, yielding performance gains of approximately 2–9 percentage points across datasets and evaluation metrics.
+This indicates that classic hand-defined features are too simple for this task and highlights the potential of using deep embeddings for data-limited applications such as NFI classification.
+By leveraging openly available satellite data and deep embeddings from pre-trained models, this approach significantly improves classification accuracy compared to traditional methods and can effectively complement existing forest inventory processes.
 keywords:
-National Forest Inventory , deep learning , pre-trained remote sensing foundation model , time series classification
+National Forest Inventory , deep embeddings , pre-trained remote sensing model , time series classification
 journal: International Journal of Applied Earth Observation and Geoinformation
 […]
-[first]organization=Wageningen University,addressline=Droevendaalsesteeg 3 , city=Wageninegn, postcode=6708 PB, country=The Netherlands
+[first]organization=Wageningen University,addressline=Droevendaalsesteeg 3 , city=Wageningen, postcode=6708 PB, country=The Netherlands
 […]
@@
 [third]organization=Wageningen Environmental Research,addressline=P.O. Box 47, city=Wageningen, postcode=6700 AA, country=The Netherlands
+[…]
+[fourth]organization=University of Bonn,addressline=Meckenheimer Allee 174, city=Bonn, postcode=53115, country=Germany
 1 Introduction
-Forests play a significant role in mitigating climate change, adopting disaster prevention strategies, and providing ecosystem services, including sequestering carbon dioxide (CO2), providing wood materials, and serving as a source of biodiversity (Tomppo et al., 2010; FAO, 2020; Francini et al., 2024).
-Tree species diversity improves productivity of materials and resistance to natural disturbance (Jactel et al., 2017).
-Therefore, monitoring and tracking records of forests including spatial distribution of tree species are necessary for a sustainable forest management.
-National Forest Inventory (NFI)s are the primary source of information for various purposes such as sustainable forest management, industry investment planning, biodiversity monitoring, and Greenhouse Gas accounting (Tomppo et al., 2010; Bonannella, 2024).
-The Paris Agreement also requires the submission of a national carbon inventory, including carbon removals and reductions from forest lands (UNFCCC, 2015).
+Forests provide critical climate-mitigation and ecosystem services, and tree-species diversity increases productivity and resistance to disturbances (Tomppo et al., 2010; FAO, 2020; Francini et al., 2024; Jactel et al., 2017).
+As a result, timely spatially explicit information on tree-species distribution is increasingly needed to support sustainable forest management, biodiversity assessments and carbon accounting (Tomppo et al., 2010; Bonannella, 2024; UNFCCC, 2015; Schelhaas et al., 2014; Hermosilla et al., 2022; Blickensdörfer et al., 2024).
+National Forest Inventory (NFI)s are the primary national reference for these applications, but they are fundamentally constrained by sampling-based field measurements typically repeated every 5–10 years, limiting their ability to capture rapid climate- and land-use-driven change at relevant scales (Tomppo et al., 2010; Bonannella, 2024).
 Spatial tree species distribution information plays an important role in NFIs for various applications such as carbon storage estimation, forest management, and biodiversity assessments (Hermosilla et al., 2022; Blickensdörfer et al., 2024).
 Additionally, detailed tree species information is essential for national reports to the Forest Resource Assessment (FRA) of the Food and Agriculture Organization (FAO) and Forest Europe (Schelhaas et al., 2014).
-One of the key challenges in managing forest inventory is frequency and scale.
-Traditional inventory methods are based on sample-based field measurements conducted every 5 to 10 years (Tomppo et al., 2010).
-Current climate change and land-use changes due to economic growth cannot be captured in a timely manner using these traditional methods (Bonannella, 2024).
-Remote sensing helps improve and update NFIs (Hermosilla et al., 2022; Francini et al., 2024).
-Satellite sensors capture information about Earth every few days to several weeks, allowing us to monitor Earth’s condition more frequently and extensively, even in remote areas.
-However, the global coverage and complex data volumes in remote sensing make human visual interpretation impractical for large-scale analysis.
-Machine learning techniques have achieved significant improvements in Earth observations, including tree species classification, particularly when using multiple timestamps of images over large areas (Blickensdörfer et al., 2024; Francini et al., 2024; Hermosilla et al., 2022).
-The RF algorithm is one of the most popular machine learning algorithms for tree species classification due to its robustness, interpretability, and ability to handle high-dimensional data (Breiman, 2001).
-However, RF models require well-designed input features engineering, and this selection of appropriate features is crucial for model performance (Heaton, 2016).
+Remote sensing has therefore become a key component of enhanced NFIs providing frequent, wall-to-wall observations that can complement plot networks and enable more up-to-date mapping (Hermosilla et al., 2022; Francini et al., 2024; White et al., 2025).
+Yet the scale and complexity of satellite time series make manual interpretation infeasible, motivating automated approaches.
+Machine learning methods, especially those leveraging multi-temporal imagery, have improved performance in Earth observation applications including tree-species classification (Blickensdörfer et al., 2024; Francini et al., 2024; Hermosilla et al., 2022).
+In operational NFI workflows, this has largely been addressed with Random Forest (RF) models over engineered time-series predictors, which perform well but hinge on careful feature design and local domain knowledge.
+The Random Forest algorithm is one of the most popular machine learning models for tree species classification due to its robustness, interpretability, and ability to handle high-dimensional data (Breiman, 2001).
+However, RF models require well-designed feature engineering, and the selection of appropriate features is crucial for model performance (Heaton, 2016).
 These choices depend on domain knowledge and target area characteristics such as climate and tree species variety, and often fail to include all necessary features (Ahlswede et al., 2023).
-In the context of NFIs, several state-of-the-art machine learning models utilizing RF have emerged (Hermosilla et al., 2022; Blickensdörfer et al., 2024; Francini et al., 2024)).
-These current methods for tree species classification rely on country-specific knowledge for input features, parameter settings, and high-quality data.
-However, scaling these methods to other countries at a national level presents challenges due to cost constraints and data availability limitations.
-Furthermore, deep learning models, which generally require large datasets to enhance performance, have remained largely unexplored for national-level tree species classification, primarily due to the limited size of available training datasets.
-These constraints collectively hinder the transferability of existing methods across different regions or countries.
-Deep learning models such as transformer architectures (Vaswani et al., 2023) have been recently introduced for forest monitoring, including tree species classification.
+In the context of NFIs, several state-of-the-art machine learning models utilizing RF have emerged (Hermosilla et al., 2022; Blickensdörfer et al., 2024; Francini et al., 2024)).The recent models have taken advantage of Satellite Image Time Series (SITS) to capture phenology differences among tree species (Hemmerling et al., 2021; Grabska and Sulikowska, 2025).
+These current methods for tree species classification rely on country-specific knowledge for selection and calculation of input features.
+Even though those models have achieved promising results, generalization and improvement of input data for model training remains the question.
+Deep learning models such as transformer architectures (Vaswani et al., 2023) have been recently introduced for forest monitoring, including tree species classification (Ma et al., 2024).
 This adoption is driven by increasing interest in multimodal and time-series data fusion in Remote Sensing (RS), enabled by the availability of big data and advancements in deep learning models (Li et al., 2022).
-Deep learning models can capture complex patterns in input data and create deep features that can be used for downstream tasks with classifiers and regressors including RF (Basu et al., 2015).
+Earth embeddings (Klemmer et al., 2025) from pre-computed features of deep learning models can capture complex patterns in in an easy-to-access form that can be directly used for downstream tasks with classifiers and regressors including RF (Basu et al., 2015).
+It is evident that models such as the temporal transformer have been introduced to capture complex temporal dependencies (Rußwurm_multi-temporal-land-classification_2018; Sainte Fare Garnot et al., 2020).
 While recent studies have achieved success in regional-scale tree species classification using high-quality labeled data, significant gaps remain in large-scale classifications (Fassnacht et al., 2016) for NFIs due to limited labeled data availability and high computational cost for training.
-Freely available pre-trained models, trained on large unlabeled datasets containing millions of pixels or images, have emerged as powerful tools for various downstream tasks.
-These models can achieve comparable or superior accuracy to traditional state-of-the-art machine learning approaches through fine-tuning without computationally expensive pre-training (Bommasani et al., 2022).
+Freely available pre-trained models and derived embeddings, trained on large unlabeled datasets containing millions of pixels or images, have emerged as powerful tools for various downstream tasks (Tseng et al., 2024; Feng et al., 2025; Brown et al., 2025).
+These models and embeddings can achieve comparable or superior accuracy to traditional state-of-the-art machine learning approaches through fine-tuning without computationally expensive pre-training (Bommasani et al., 2022).
 Self-supervised learning, where models are trained without labels, has gained particular attention in RS applications (Wang et al., 2022).
@@
 The rapid growth of both labeled and unlabeled datasets for RS (Gorelick et al., 2017; Ahlswede et al., 2023) has enabled the development of various pre-trained models for tasks including tree classification (Lu et al., 2024).
-However, research comparing performance between fine-tuned pre-trained models and traditional approaches for tree species classification in NFIs remains limited.
-This paper evaluates the effectiveness of deep features from freely available fine-tuned pre-trained models compared to a recently published random forest approach by Francini et al.
-(2024), who set a new state-of-the-art for Tree Species Identification in the Netherlands through a RF fitted on median reflectances and harmonic features.
-The paper is structured as follows: section˜2 details the three datasets from the Dutch Forest Inventory used in this study.
-Section˜3 details the deep features we fine-tuned together with the harmonic features, as proposed by prior work.
+Despite rapid progress in machine learning for remote sensing, a key scientific challenge remains: under limited labeled data, can deep feature embedding representations capture spatial and temporal detail required by national forest inventories better than existing hand-crafted models.
+While existing approaches, such as country-specific RF models, depend heavily on expert-designed features and do not easily transfer to other ecological or sensing conditionsnew pre-trained embedding models promise scalable feature extraction but have not yet been systematically assessed for operational forest inventory settings.
+In short, this study asks whether representations learned from large, unlabeled remote sensing datasets can outperform or complement traditional hand-crafted features in NFI workflows.
+Our main contributions are threefold:
+  1. 1.
+    We identify the lack of transferable, data-efficient tree species mapping methods as a limiting factor for timely and scalable forest monitoring.
+  2. 2.
+    We integrate deep embeddings from fine-tuned self-supervised models into a tree species classification pipeline, directly comparable to a state-of-the-art RF approach (Francini et al., 2024).
+  3. 3.
+    Using Dutch National Forest Inventory data, we conduct a systematic comparison between harmonic + medoid features and fine-tuned deep embeddings representations, analyzing accuracy, robustness, and generalization potential.
+2 Methods
+In this study, we fundamentally study the impact of time series feature embeddings on national forest inventory (NFI) by comparing hand-designed features from Francini et al.
+(2024), which represent the current state-of-the-art for Dutch NFI mapping, with deep learning embeddings from Presto, Tessera, and AlphaEarth.
+Table˜1 provides a high-level comparison in terms of data and usability criteria and the following-subsections summarize the respective models.
+Time series input is relevant to capture phenological sequential patterns, downloadable embeddings make applying deep embeddingsvery user-friendly and applicable for large-scale mapping tasks, and an open-source available fine-tunable model allow further modifications and possibly better downstream accuracy.
+Hence, we outline the methodological details of both complementary approaches in the next two sections before detailing the experimental setup in section˜4.
+Table 1: Comparison of deep learning embedding models Presto (Tseng et al., 2024), TESSERA (Feng et al., 2025), and AlphaEarth (Brown et al., 2025) with traditional features (Francini et al., 2024) across time data and usability criteria relevant to NFI applications.
+Additional information includes encoder parameter count, pretraining dataset size, and pretraining type.
+Model Time Series Input Open
+Embeddings
+Open
+Model
+Encoder Parameters Pretraining Data Size Pretraining Type
+Presto (Tseng et al., 2024) 12 fixed monthly steps No Yes 0.4M 21.5M pixels Pixel-level
+TESSERA (Feng et al., 2025) 40 random steps/year Yes (on request) Yes (on request) 40M 0.8B pixels Pixel-level
+AlphaEarth (Brown et al., 2025) No (implicit temporal summarization) Yes No 480M–1B 3B images Image-level
+Harmonic + Medoid features (Francini et al., 2024) Harmonic coefficients and seasonal median No N/A N/A N/A N/A
+2.1 Deep Embeddings
+These models reflect a recent development toward making pre-trained deep embeddings openly available (open embeddings) either directly for download, or through open-source, pre-trained, and fine-tunable models (open model).
+2.1.1 Presto “Pretrained Remote Sensing Transformer” (Tseng et al., 2024)
+Architecture and Training: Presto is a lightweight pixel-based Transformer trained on explicit multi-sensor time series.
+Each pixel is represented by a 12-month sequence of Sentinel-1 (S-1), Sentinel-2 (S-2), NDVI, ERA5 climate, and Dynamic World land-cover data, augmented with static contextual variables such as SRTM elevation and geographic coordinates.
+Presto learns through self-supervised masked autoencoding, reconstructing missing timesteps and sensor channels to capture temporal dependencies and handle missing data.
+Suitability for NFI: Because Presto directly models monthly vegetation trajectories and integrates environmental context, it is highly suited for national forest inventory (NFI) tasks where phenological signals (e.g., seasonal greenness, moisture, canopy change) dominate.
+It can characterize both intra-annual and inter-annual forest dynamics.
+The trade-off is that Presto must be run locally, requiring access to raw satellite data, but it offers the strongest temporal interpretability for NFI-type monitoring.
+Accessibility of model and embeddings: The Presto model is openly available (e.g., via GitHub) for download and local execution.
+However, pre-computed embeddings are not published for direct download — users must compute embeddings themselves from raw time-series inputs.
+This implies a heavier setup but full control over fine-tuning and feature extraction.
+We detail our setup in the section˜A.1.
+2.1.2 Tessera “Temporal Embeddings of Surface Spectra” (Feng et al., 2025)
+Architecture and Training: Tessera uses a dual-encoder Transformer architecture—one branch for S-1 SAR and another for S-2 optical data—processing entire annual time series per pixel (timestep $\times$ channels).
+Its self-supervised pretraining applies a modified Barlow Twins loss that enforces invariance between partial temporal views of the same pixel, encouraging the model to learn phenologically stable annual representations.
+The output is a 128-dimensional annual embedding per 10 m pixel, globally covering 2017–2024, available as pre-computed data on request.
+Suitability for NFI: Tessera is well suited for NFI applications because it directly encodes annual spectral and backscatter evolution—the key seasonal information driving forest structure and productivity estimates—while being readily accessible as pre-computed embeddings.
+However, since it outputs annual summaries rather than continuous time series, it provides a coarser representation of intra-annual phenology compared to Presto.
+Accessibility of model and embeddings: The Tessera repository is publicly accessible (MIT license) and supports inference and embedding extraction.
+Pre-computed embeddings are provided via the accompanying GeoTessera library for many years and regions, subject to request.
+Fine-tuning the model is possible since the code and architecture are open, but the underlying model weights need to be requested.
+2.1.3 AlphaEarth Foundations “Embedding Field Model” (Brown et al., 2025)
+Architecture and Training: AlphaEarth integrates diverse Earth observation sources (S-1/2, Landsat, GEDI, ERA5-Land, GRACE, and others) through a Space-Time Precision (STP) encoder that combines spatial and temporal attention with convolutional operations.
+It is trained as a multi-modal autoencoder with implicit decoders that reconstruct missing observations across time and modalities, regularized by a teacher–student consistency and contrastive alignment loss.
+The model learns continuous-time latent fields, producing annual “embedding layers” distributed globally via Google Earth Engine (Gorelick et al., 2017).
+Suitability for NFI: While AlphaEarth embeddings reflect long-term temporal consistency and cross-sensor coherence, they do not directly input dense time series per pixel.
+Instead, they summarize information over broader “valid periods,” emphasizing spatial and modal fusion rather than explicit phenological sequencing.
+This makes AlphaEarth moderately suited for NFI applications—excellent for stable, multi-sensor mapping of forest attributes, but less precise for capturing fine-scale seasonal forest dynamics critical to phenology-based inventories.
+Accessibility of model and embeddings: The embedding dataset of AlphaEarth (annual 10 m resolution embeddings) is openly accessible through the Earth Engine Data Catalog.
+However, the full trained model weights (for fine-tuning or re-training) are not officially released in full form by the authors.
+Therefore, end users can readily apply the embeddings, but cannot fine-tune the original model out of the box.
+2.2 Harmonic and Seasonal Features
+We reproduced closely the feature engineering approach of Francini et al.
+(2024), who published a new state-of-the-art classification approach for tree species mapping in the Netherlands.
+Their approach generate 209 features containing Seasonal Medoid Statistics and Harmonic Features that we describe in the next two paragraphs.
+While Francini et al.
+(2024) used only features extracted from S-2 time series, we found experimentally (detailed later in section˜5.1 and table˜3) that including features from S-1 further slightly improves the results.
+Since we will also use S-1 to extract deep embeddings (section˜2.1), we include S-1 features to ensure a fair comparison.
+76 Seasonal Medoid Statistics
+In total, 12 bands from S-1 and S-2 are selected and additional 7 indices are calculated from S-2 including Normalized Difference Vegetation Index (NDVI), Normalized Burn Ratio (NBR), Enhanced Vegetation Index (EVI), and Tasseled Cap transformations: Brightness (TCB), Wetness (TCW), Greenness (TCG), Angle (TCA).
+This results in 19 bands, where for each band the median band value in each of the 4 seasons, winter (January, February, December), spring (March-May), summer (June-August), and autumn (September-November), the medoids values are extracted.
+These features describe the data through season-wise statistics, but do not capture the dynamic change in seasonality itself.
+133 Harmonic Features
+To capture seasonal changes, sine and cosine harmonics
+$P_{t}=\beta_{0}+\beta_{1}t+\beta_{2}\cos(2\pi\omega t)+\beta_{3}\sin(2\pi\omega t)$ (1)
+are fitted to the underlying monthly signal of each band according.
+The seven harmonic parameters $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $A$, $\phi$ for each of the 19 band are used as classification features alongside the residual RMSE: $\beta_{0}$ is the constant, $\beta_{1}$ is the time coefficient, and $\beta_{2}$ and $\beta_{3}$ are the frequency sine and cosine coefficients, respectively.
+To fit these four coefficients and to select $P_{t}$, the pixel $p$ harmonic values at time $t$, we used a least squares regression to fit Eq.
+1.
+$A=\sqrt{\beta_{2}^{2}+\beta_{3}^{2}}$ is the amplitude of the harmonic curve on the y-axis, $\phi=\arctan(\beta_{3}/\beta_{2})$ is the phase of the curve on the x-axis to the origin and $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(P_{t}-X_{t})^{2}}$ is root mean square error between $P_{t}$ and the actual pixel values $X_{t}$.
+The frequency $\omega$ = 1 (in years) and indicates one cycle per unit of time, initial guess values before fitting four coefficients were set as $\beta_{0}=0.1$, $\beta_{1}=0.1$, $\beta_{2}=0.4$, $\beta_{3}=0.4$.
+This approach was proposed by Francini et al. (2024) based on S-2 features. We added of S-1 features, as we found them to be slightly beneficial in accuracy (see section˜5.1).
 3 Data
-We use three datasets from the Dutch National Forest Inventory (NFI) at different complexity levels that we will use to compare the proposed fine-tuned deep time series features with the RF by Francini et al.
-(2024) fitted on harmonic time series features, as detailled later in section˜3.
-Complex & imbalanced (COMB)
-13 tree species classes, imbalanced distribution, 1,462 time series samples.
-Simplified & imbalanced (SIMB)
-7 tree species classes, imbalanced class distribution, 1,479 time series samples.
-Simplified & balanced (SIBA)
-7 species classes, a balanced class distribution, by Francini et al. (2024). In total 13,790 samples and contains additional samples augmented by photointerpretation.
-Table˜1 shows a detailed list of the tree species across all three datasets.
 Table 2: Species count and grouping in NFI and Francini datasets.
@@
 3.2 National Forest Inventory
+We use three datasets from the Dutch National Forest Inventory (NFI) at different complexity levels that we will use to compare the proposed deep embeddings with the RF by Francini et al.
+(2024) fitted on harmonic time series features, as detailed later in section˜2.
+Despite the relatively small overall forest area in the Netherlands (approximately 10–11 % of land cover), forests display spatial variability in species composition, age structure, and mixed stand types, reflecting responses to site conditions, management history, and ecological processes (Schelhaas et al., 2014).
+Complex & imbalanced (COMB)
+13 tree species classes, imbalanced distribution, 1,462 time series samples.
+Simplified & imbalanced (SIMB)
+7 tree species classes, imbalanced class distribution, 1,479 time series samples.
+Simplified & balanced (SIBA)
+7 species classes, a balanced class distribution, as proposed by Francini et al.
+(2024).
+In total 13,790 samples and contains additional samples augmented by photointerpretation from Dutch NFI samples.
+Orginal Dutch NFI samples (COMB and SIMB) are derived by random sampling with a density of 1 point per 100 ha(Schelhaas et al., 2014), which means that each random point is designated per square kilometer, and SIBA removed adjacent pixels with minimum distance among sample pixels was set to 15m, which reduced issues associated with spatial autocorrelation(Francini et al., 2024; Hermosilla et al., 2022).
+Table˜2 shows a detailed list of the tree species across all three datasets.
 Label Data Origin Ground truth data were collected manually through field measurements at 3,062 plots for the Dutch National Forest Inventory 6 between 2012 and 2013.
 Each plot contained a circular area with a variable radius (5 to 20 m) to ensure inclusion of at least 20 trees (Schelhaas et al., 2022).
-These plots correspond to 1 to 16 pixels at a $10\times 10$m resolution of Sentinel-1 and Sentinel-2.
+These plots correspond to 1 to 16 pixels at a $10\times 10$m resolution of S-1 and S-2.
 Due to privacy considerations, the precise coordinates of the plots were obtained under a confidentiality agreement with the Dutch government, with the requirement that the data be discarded after project completion.
@@
 Class Aggregation The original NFI data contains 19 dominant tree species classes.
-However, we noticed that six dominant species classes have less than 10 samples and we decided to remove these classes to not extenstively rely in our evaluation on underrepresented tree species, similar to Kang et al.
+However, we noticed that six dominant species classes have fewer than 10 samples and we decided to remove these classes to not extensively rely on underrepresented tree species in our evaluation, similar to Kang et al.
 (2017).
@@
 We call Simplified & Balanced (SIBA) compared to aforementioned ones, as it contains 7 tree species classes and an even number of samples per class.
-2.3 Remote Sensing Time Series
-Since mono-temporal data cannot adequately capture seasonal leaf phenology for tree species classification, we utilized time-series data spanning January through December 2020.
+To augment the original Dutch NFI plots, the authors performed a rigorous three-step validation process: (1) visual photointerpretation using high-resolution orthophotos and Google satellite imagery to remove plots affected by land-cover change or mixed-species compositions; (2) stand-polygon identification to extract homogenous spectral samples; and (3) spatial thinning, where a minimum distance of 15 m was maintained between pixels to minimize spatial autocorrelation and label noise.
+The final dataset comprises 13,790 samples (1,970 per class).
+3.3 Remote Sensing Time Series for Presto and harmonic + medoid calculation
+We extracted time-series data spanning January through December 2020 through Google Earth Engine (Gorelick et al., 2017).
 This temporal range aligns with both the Pretrained Remote Sensing Transformer (Presto) model requirements (Tseng et al., 2024) and previous Dutch tree species classification research (Francini et al., 2024).
-Input Data Modalities S2, S1, ERA5, SRTM.
+Input Data Modalities S-2, S-1, ERA5, SRTM.
 This study employs a comprehensive multi-source remote sensing (RS) approach that integrates satellite imagery and environmental data to characterize forest vegetation across space and time.
-The model relies on several key data sources: Sentinel-1 SAR GRD data with VH and VV polarizations at a 6-day revisit frequency; Sentinel-2 multispectral Level 1C data with 5–10 day revisit frequency, including 10 spectral bands (B2–B8, B8A, B11, B12) and NDVI computed during preprocessing; ERA5 climate data providing monthly measurements of 2m air temperature and total precipitation; and static SRTM terrain data comprising elevation and slope.
-Remote Sensing Data Preparation For S2 data, cloud masking was performed using the S2 cloud probability dataset (Pasquarella et al., 2023), excluding pixels with cloud cover probability exceeding 65% (Francini et al., 2024).
+The model relies on several key data sources: S-1 SAR GRD data with VH and VV polarizations at a 6-day revisit frequency; S-2 multispectral Level 1C data with 5–10 day revisit frequency, including 10 spectral bands (B2–B8, B8A, B11, B12) and NDVI computed during preprocessing; ERA5 climate data providing monthly measurements of 2m air temperature and total precipitation; and static SRTM terrain data comprising elevation and slope.
+Using the plot-level coordinates, we extracted a single set of time-series data (or embeddings) for each plot; consequently, the COMB dataset comprises 13,790 plots and 13,790 corresponding samples.
+During preprocessing, satellite data with coarser spatial resolutions than 10 m (i.e., S-1, ERA5, and SRTM) were resampled to a common spatial resolution of 10 m to ensure consistency with Presto input(Tseng et al., 2024) and harmonic + medoid feature input(Francini et al., 2024).
+Remote Sensing Data Preparation For S-2 data, cloud masking was performed using the S-2 cloud probability dataset (Pasquarella et al., 2023), excluding pixels with cloud cover probability exceeding 65% (Francini et al., 2024).
 Geographic coordinates were utilized both for satellite data extraction from the GEE archive and as input features for the model.
-Raw time-series data was downloaded as individual CSV files for each plot location across all datasets (NFI, Francini, and non-labeled data), ensuring direct correspondence between remote sensing observations and ground truth measurements.
-We calculated monthly median values for each band in S1 and S2 except monthly ERA5 data.
-After getting monthly median values for S1 and S2, Normalized Difference Vegetation Index (NDVI) was calculated from S2 bands to capture vegetation dynamics.
-3 Methods
-In this study, we fundamentally compare results between a classic machine learning pipeline using a RF fitted on harmonic features and seasonal reflectances by Francini et al.
-(2024) with pre-trained deep-learning features obtained by fine-tuning the Presto time series foundation model proposed by Tseng et al.
+Raw time-series data was downloaded as individual CSV files for each plot location across all datasets (SIMB, COMB and SIBA), ensuring direct correspondence between remote sensing observations and ground truth measurements.
+4 Experimental Setup
+This study compares classic machine learning with hand-crafted feature engineering against deep learning approaches for tree species classification.
+Our experimental design follows a systematic progression to evaluate different methodological approaches and their combinations.
+4.1 Comparative evaluation framework
+Our experimental design implements four distinct comparison scenarios to systematically evaluate feature representations and classifier architectures:
+  1. 1.
+    Sensor contribution analysis: S-1 vs S-2 vs S-1+S-2 using harmonic + medoid features with RF
+  2. 2.
+    Feature representation comparison: Harmonic + medoid features vs deep embeddings using RF
+  3. 3.
+    Classifier architecture evaluation: RF vs Multi Layer Perceptron (MLP) using deep embeddings
+This systematic approach allows us to isolate the contributions of different methodological components while building toward the optimal configuration for operational NFI applications, extending the work of Francini et al.
+(2024) with comprehensive deep learning comparisons.
+4.2 Model fine-tuning
+The fine-tuning approach leverages the pre-trained Presto encoder’s learned representations while adapting the model to the specific characteristics of Dutch forest species.
+This transfer learning strategy balances computational efficiency with task-specific optimization.
+For fine-tuning the pre-trained models, the encoder block from the pre-trained model was reused and trained with newly attached Multi Layer Perceptron (MLP) classifier for the downstream task.
+This fine-tuning model is trained on the 1,479 and 13,790 samples for 7 aggregated groups classification task (SIMB, SIBA), and the 1,462 samples for the 13 dominant species classification task (COMB).
+Pre-trained encoder of Presto was set as trainable in this process.
+Based on the current state-of-the-art classifier for tree species classification (Mouret et al., 2025), the MLP classifier architecture consists of a 3-layer MLP with no dropout (rate: 0.0).
+The layer configuration comprises 1024 nodes in the first layer, 512 nodes in the second layer, and 256 nodes in the final layer, with batch normalization and ReLU activation applied throughout.
+Training parameters are configured as follows: a learning rate of 0.0001 with weight decay of 0.00746.
+The model is trained for a maximum of 100 epochs using a batch size of 64, with cross-entropy loss for optimization.
+The best fine-tuned model and its accuracy were selected based on validation loss.
+All fine‑tuning and embedding extraction experiments were conducted on Google Colab using an NVIDIA Tesla T4 GPU (16GB GPU memory).
+Each fine‑tuning run used a maximum of 100 training epochs, but due to early stopping based on validation performance, most runs terminated in fewer than 30 epochs in SIMB (1,479 samples) and COMB (1,462 samples), 50 epochs in SIBA (13,790 samples).
+As a result, a single fine‑tuning run on the smaller SIMB and COMB datasets takes approximately 1 minute, with embedding extraction requiring an additional  1 minute per dataset.
+For the larger SIBA dataset, a single fine‑tuning run takes approximately 10 minutes and embedding extraction takes  3 minutes.
+4.3 Data splitting strategy
+Out-of-bag (OOB) error estimate with RF algorithm was applied in Francini et al.
+(2024), however we also need to evaluate MLP classifier in a comparable setting with RF classifier.
+Based on prior research (Blickensdörfer et al., 2024; Freudenberg et al., 2025b; Hermosilla et al., 2022), a 70/30 (approximately 2:1 ratio) train-test split was applied for fine-tuning and RF classification.
+In addition, for MLP, training data is split into MLP train data and MLP validation data with a ratio of 80:20 (56%, 14% of the total dataset) to select the best model by lowest validation loss, which ensures that test data is not used during training and reduces overfitting.
+During fine-tuning and classification, the same training and test data were used across datasets and models.
+This practice ensures that comparable results across the model with same train-test split and extracted embeddings are not trained on test data.
+By adopting stratified sampling, each class is equally split into train (MLP train data and MLP validation data) and test data at this ratio.
+Appendix table A.1 shows species class distribution across train (MLP train and MLP validation) and test splits for three datasets.
+As described in Section˜3, our datasets have already mitigated spatial autocorrelation.
+In addition, our datasets especially NFI data (SIMB and COMB) are small (1,462 to 1,479) and have a lot of classes (6 to 13 classes), which makes spatial aware split with class-wise stratified sample difficult.
+Therefore, we apply stratified random splits.
+We repeat each experiment five times, applying different random seeds to generate independent data splits.
+.
+4.4 Training pipeline
+In each random seed, we split each dataset for 70% training (56% MLP training, 14% MLP validation) and 30% test.
+MLP training dataset is used for fine-tuning Presto decoder with MLP classifier, and MLP validation data is used to decide the best model.
+Using this best fine-tune Presto model with MLP classifier, we evaluate test data (MLP fine-tune Presto result).
+In addition, by using fine-tune Presto decoder, we extracted embeddings for all samples in the dataset.
+Those fine-tune presto embeddings as well as other embeddings (AEF and TESSERA) are splitted in the exact same manner as MLP fine-tuning.
+70% of those embeddings are used for RF training and 30% are used for RF evaluation.
+4.5 Random Forest classifier
+Following the experimental setup of Francini et al.
+(2024), we implemented a Random Forest (RF) classifier (Breiman, 2001) consisting of 500 trees.
+The number of features considered at each split was set to the square root of the total number of input predictors.
+4.6 Performance evaluation
+Model performance was evaluated using a comprehensive set of established accuracy metrics: per-class recall, F1 scores (Macro, Weighted), confusion matrices, and overall accuracy.
+While overall accuracy provides a general assessment of classification performance, it exhibits reduced sensitivity to class imbalance issues that are prevalent in ecological datasets.
+Therefore, the F1 score (Macro, Weighted), which represents the harmonic mean of precision and recall, was employed as a supplementary metric to provide a more balanced evaluation of model performance when addressing the inherent class imbalances in tree species distribution data.
+Confusion matrices were used to visualize the model’s performance at the class level.
+These metrics are widely used in tree species classification studies (Goutte and Gaussier, 2005; Hermosilla et al., 2022; Francini et al., 2024; Blickensdörfer et al., 2024) and provide comprehensive assessment of classification performance across all species classes.
+4.7 Statistical validation
+To evaluate variability of model performance, each model was evaluated five times with different random seeds, which affect to train-test split and model initialization.
+The mean and standard deviation of the accuracy metrics were calculated for each model.
+5 Results
+5.1 Influence of multi-source input features in harmonic and medoid features
+Following the approach of Francini et al.
+(2024), we first evaluated the contribution of S-1 radar data to optical S-2 classification.
+Using RF with harmonic + medoid features, the addition of S-1 provided modest improvements: 0.4% for the 13-class COMB dataset and 1.4% for the 7-class SIMB dataset.
+These results justify the inclusion of S-1 features in our subsequent comparisons.
+We first evaluated whether combining optical and radar satellite data improves tree species classification using hand-crafted features in a RF classifier.
+Feature sets were derived from seasonal medoid composites and harmonic fits, using either S-2 alone or in combination with S-1.
+The goal was to assess whether radar backscatter, which captures structural and moisture-related canopy properties, offers complementary information to optical reflectance data.
+As shown in Table 3 the inclusion of S-1 provided modest but consistent improvements across both datasets.
+For the 13-class dataset (COMB), accuracy increased from 66.34% to 66.71% when adding S-1 to S-2 derived features.
+In the 7-class dataset (SIMB), accuracy improved from 72.87% to 74.22%.
+These results suggest that the radar signal contributes marginally to distinguishing between species, especially under a more simplified class grouping.
+While the absolute gains remain small, incorporating S-1 ensures a more robust feature representation, particularly in settings where optical signals may be limited by cloud cover or seasonal variability.
+The modest accuracy gains observed here are also useful to contextualize the more substantial differences in performance reported in the next sections, which stem from the feature representations rather than the data sources themselves.
+Table 3: Classification accuracy comparison across different feature combinations and satellite data sources in NFI data (in %). Bold values indicate best performance for each label type.
+Label Features Sentinel-1 Sentinel-2 S-1 + S-2 Combined
+Band Acc. Band Acc. Band Acc.
+COMB (13 classes) Seasonal (S) S-1 51.03 S-2 64.03 S-1 + S-2 64.52
+Harmonic (H) S-1 55.04 S-2 63.79 S-1 + S-2 64.64
+All (S+H) S-1 57.72 S-2 66.34 S-1 + S-2 66.71
+SIMB (7 classes) Seasonal (S) S-1 55.60 S-2 70.18 S-1 + S-2 71.93
+Harmonic (H) S-1 60.19 S-2 71.39 S-1 + S-2 71.52
+All (S+H) S-1 62.75 S-2 72.87 S-1 + S-2 74.22
+5.2 Deep Embeddings vs Harmonic features
+We compared classification performance between traditional hand-crafted features (harmonic and medoid features) and several deep embeddings as described in table 4.
+We first analyze the effect of different feature representations while fixing the classifier to Random Forest (RF), allowing a direct comparison between deep embeddings and traditional harmonic + medoid features, as well as between frozen and fine-tuned Presto representations.
+Across all datasets, RF models trained on deep embeddings consistently outperform those using harmonic + medoid features.
+In particular, results obtained from embeddings with the RF classifier show that fine-tune Presto embeddings consistently outperform frozen Presto, AEF, and TESSERA embeddings across all datasets and evaluation metrics, including overall accuracy, F1 macro, and F1 weighted scores.
+This highlights the benefit of task-specific fine-tuning for adapting general-purpose deep embeddings to forest species classification.
+On the SIMB dataset, which is small and moderately imbalanced, RF with fine-tuned Presto embeddings achieved an overall accuracy of 73.02% (±1.55), an F1 macro score of 57.60% (±1.74), and an F1 weighted score of 71.80% (±1.69), compared to 69.48% (±0.73) accuracy, 51.03% (±1.71) F1 macro, and 66.98% (±1.06) F1 weighted for harmonic + medoid features.
+This corresponds to gains of approximately 3.5 percentage points in accuracy, over 6 points in F1 macro, and nearly 5 points in F1 weighted, highlighting the advantage of learned representations in low-sample regimes while accounting for class frequency For the more challenging COMB dataset, characterized by a larger number of classes and stronger class imbalance, the performance gap remains evident.
+RF with fine-tuned Presto embeddings reached 65.91% (±1.72) overall accuracy, 46.11% (±3.00) F1 macro, and 62.34% (±1.97) F1 weighted, compared to 62.80% (±0.91) accuracy, 38.95% (±2.02) F1 macro, and 57.13% (±1.25) F1 weighted using harmonic + medoid features.
+The substantial improvement of 3 points in accuracy, 7 points in F1 macro and 5 points in F1 weighted indicates that deep embeddings better capture discriminative information for minority classes under increased class complexity, while gains in F1 weighted reflect consistent improvements across dominant classes.
+The largest absolute performance differences are observed on the SIBA dataset, which is both balanced and substantially larger.
+Here, RF with fine-tuned Presto embeddings achieved 93.37% (±0.69) accuracy, 93.33% (±0.69) F1 macro, and 93.33% (±0.69) F1 weighted, compared to 84.47% (±0.41) accuracy, 84.40% (±0.40) F1 macro, and 84.40% (±0.40) F1 weighted for harmonic + medoid features.
+The gains of nearly 9 points across evaluation matrices reflect the balanced class distribution of SIBA and demonstrate that deep embeddings scale effectively with data availability and clearer class separation.
+Within RF-based deep models, fine-tuning Presto consistently improves performance over frozen embeddings, particularly for SIMB and SIBA.
+The gains are more modest for COMB, suggesting that while fine-tuning enhances representation quality, class imbalance and higher label complexity still limit achievable performance.
+We next examine the influence of classifier architecture by comparing the MLP classifier for tree species classification Mouret et al.
+(2025) with an RF classifier, both trained on the same deep embeddings extracted from the fine-tuned Presto model.
+This isolates the effect of the classifier while keeping the input representation fixed.
+Across all datasets, the MLP classifier slightly outperforms RF when using fine-tuned Presto embeddings, with the magnitude of the performance differences varying according to dataset size and class balance.
+On the SIMB dataset, MLP achieved 73.67% (±1.29) accuracy, 62.29% (±4.60) F1 macro, and 73.05% (±1.52) F1 weighted, compared to RF’s 73.02% (±1.55) accuracy, 57.60% (±1.74) F1 macro, and 71.80% (±1.69) F1 weighted.
+While overall accuracy is comparable and F1 macro exhibits higher variance due to the small and imbalanced nature of SIMB, MLP attains marginally higher F1 macro and weighted scores than RF, indicating slightly improved average performance across classes and stronger performance on majority classes, resulting in better overall balance when accounting for class frequency.
+For the COMB dataset, the advantage of MLP becomes more pronounced.
+MLP achieved 67.40% (±1.47) accuracy, 50.76% (±2.23) F1 macro, and 65.51% (±1.29) F1 weighted, outperforming RF’s 65.91% (±1.72) accuracy, 46.11% (±3.00) F1 macro, and 62.34% (±1.97) F1 weighted.
+This suggests that the non-linear decision boundaries learned by MLPs are slightly better suited for complex, multi-class classification tasks with imbalanced distributions.
+In the SIBA dataset, both classifiers achieve strong and nearly identical performance due to the large sample size and balanced class distribution.
+MLP reached 93.54% accuracy, F1 macro, and F1 weighted, compared to 93.37% accuracy, 93.33% F1 macro, and 93.33% F1 weighted for RF.
+The negligible differences indicate diminishing returns from classifier complexity in data-rich scenarios.
+Overall, while RF offers robust and competitive performance—particularly in smaller datasets, MLPs consistently extract slightly more discriminative power from deep embeddings, especially in more complex or imbalanced classification settings, as reflected by modest improvements in both F1 macro and F1 weighted scores.
+Table 4: Classification performance comparison (in %) across feature types (deep embedding vs.
+harmonic + medoid) and classifiers (MLP vs.
+RF).
+Each model was run five times and the mean and standard deviation (±) are reported.
+Bold values indicate the best performance among MLP and RF per dataset type.
+Classification Results
+MLP Random Forest
+Dataset type Fine-tune Presto with MLP Classifier Presto
+Fine-tune Presto Harm.+med.
+(Francini et al., 2024)
+AEF
+(Frozen )
+(Brown et al., 2025)
+TESSERA
+(Feng et al., 2025)
+Overall Accuracy SIMB (7 classes) 73.67 ± 1.29 69.39 ± 2.10 73.02 ± 1.55 69.48 ± 0.73 72.74 ± 1.61 73.65 ± 1.28
+COMB (13 classes) 67.40 ± 1.47 60.68 ± 1.51 65.91 ± 1.72 62.80 ± 0.91 62.71 ± 0.56 65.28 ± 1.61
+SIBA (7 classes) 93.54 ± 1.27 84.51 ± 0.33 93.37 ± 0.69 84.47 ± 0.41 90.32 ± 0.65 75.65 ± 0.64
+F1 Macro SIMB (7 classes) 62.29 ± 4.60 47.17 ± 2.03 57.60 ± 1.74 51.03 ± 1.71 52.67 ± 2.76 57.36 ± 1.81
+COMB (13 classes) 50.76 ± 2.23 36.30 ± 2.10 46.11 ± 3.00 38.95 ± 2.02 38.05 ± 1.97 43.23 ± 3.16
+SIBA (7 classes) 93.54 ± 1.26 84.32 ± 0.36 93.33 ± 0.69 84.40 ± 0.40 90.20 ± 0.65 75.66 ± 0.62
+F1 Weighted SIMB (7 classes) 73.05 ± 1.52 66.58 ± 2.19 71.80 ± 1.69 66.98 ± 1.06 70.38 ± 1.78 71.12 ± 1.40
+COMB (13 classes) 65.51 ± 1.29 55.37 ± 1.67 62.34 ± 1.97 57.13 ± 1.25 57.18 ± 0.88 60.15 ± 1.96
+SIBA (7 classes) 93.54 ± 1.26 84.32 ± 0.36 93.33 ± 0.69 84.40 ± 0.40 90.20 ± 0.65 75.66 ± 0.62
+For direct comparison between fine-tune Presto and Harmonic + medoid features using identical S-1 + S-2 sensor inputs (detailed in Appendix Table A.2), deep embeddings show F1 score improvements of 9.0 percentage points (SIMB), 5.7 percentage points (COMB), and 7.5 percentage points (SIBA), confirming consistent performance advantages across different input configurations for embeddings.
+Appendix Table A.4 compares models using all satellite data against those using only S-1 and S-2 data for fine-tuned Presto embeddings with the MLP classifier.
+The results show no significant differences in input data for fine-tuned Presto embeddings.
+5.3 Class level performances
+To better understand class-level behavior, we visualized the confusion matrices of the best-performing model – an MLP fine-tuned on deep embeddings – for two datasets (COMB, SIBA).
+These matrices are shown in Figure 2, with each cell reporting the number of predicted samples and the corresponding recall (i.e., percentage of correct predictions among all true samples for that class).
+In the complex 13-class dataset (COMB), misclassifications are pronounced.
+Within the broadleaved group (Alnus spp., Betula spp., “Broadleaved spp.”, Fagus spp., Fraxinus spp., Populus spp., Q.
+robur/petraea, Quercus spp.), there is substantial confusion among species whose phenology or seasonal spectral signatures overlap.
+For example, Q.
+robur/petraea, being the most abundant class in this group, serves as a frequent “sink” — many other broadleaved species are misclassified as oak.
+Less common broadleaved classes, such as Alnus spp.
+and Fraxinus spp., often suffer from very poor classification performance.
+In our results, Alnus spp.
+achieves only 2.2 % recall, with 62 % of its true instances misclassified as Q.
+robur/petraea.
+Phenologically, both oak (Q.
+robur/petraea) and alder (Alnus spp.) are classified among the late-leafing deciduous species according to remote-sensing studies (Grabska-Szwagrzyk and Tymińska-Czabańska, 2024).
+In addition, species such as ash (Fraxinus spp.) are known to be difficult to discriminate reliably from other broadleaved deciduous species in remote-sensing based classification (Lisein et al., 2015).
+Because of these overlapping phenological and spectral patterns, misclassification among broadleaved species is observed.
+With conifer species (Larix spp, P.
+menziesii, P.
+sylvestris, Picea spp and Pinus spp), this group often exhibits markedly different spectral and phenological behavior compared to deciduous broadleaves, which can facilitate broad-leaf vs conifer discrimination (Vorovencii et al., 2023).
+However, misclassification still occurs within the conifer group.
+For instance, the majority of Pinus spp were misclassified as P.
+sylvestris, which is also a type of Pinus (pine).
+Thus, while the model likely performs well at the coarse level (broadleaved vs conifer), its reliability decreases when discriminating among species within each group, especially for rare species and those with phenological and spectral overlap.
+In the 7-class NFI dataset (SIMB), the model exhibits generally strong performance across most classes.
+Recall values range from approximately 90 % up to 98 % in several cases.
+For the aggregated class DarkConifer and Populus, the model yields a recall of around 98 %.
+Within the broadleaved aggregate classes, performance is more variable.
+Most of the residual confusion is observed among Quercus (Oak), Beech, and Other Broadleaf, reflecting the inherent spectral and phenological similarity within the broadleaved group.
+Figure 2: Confusion matrices for the best performing models with deep embeddings by MLP classifier.
+Complex & Imbalanced (COMB)(left), Simplified & Balanced (SIBA) (right).
+Each cell has average count and recall percentage.
+6 Discussion
+6.1 Multi-modal sensor integration for harmonic + medoid features
+Our findings on the modest but consistent benefits of combining S-1 radar with S-2 optical data for harmonic + medoid features align with recent European tree species mapping studies.
+Blickensdörfer et al.
+(2024) demonstrated that S-1 SAR data effectively fills potential data gaps in cloud-affected regions, particularly important for national-scale applications where clear-sky observation density varies substantially.
+Similarly, Liu et al.
+(2023) confirmed that S-1 provides supplementary structural information to S-2 optical data, contributing to increased reliability of large-scale tree species mapping.
+This benefit could be attributed to the complementary nature of the sensors: S-1 SAR captures structural information, while S-2 optical data provides insights into canopy density and foliage characteristics (Forkuor et al., 2020).
+Conversely, the addition of S-1 sensors for tree species classification does not always yield substantial accuracy improvements.
+Lechner et al.
+(2022) noted that the advantages of SAR are most pronounced when S-2 lacks sufficient cloud-free observations.
+Given the monthly and seasonal composite nature of our harmonic and medoid features, most S-2 composites in our study contained valid data.
+This high data availability likely resulted in the marginal improvements observed.
+Our results, showing 0.4-1.4% accuracy improvements when adding S-1 to S-2 features, corroborate these findings and provide confidence in the robustness of multi-modal approaches for operational NFI applications.
+6.2 Deep vs hand-crafted features: an emerging paradigm
+The transition from hand-crafted to learned feature representations marks a significant shift in remote sensing methodology.
+Even though the interpretability of deep features remains an active research topic (Rao et al., 2025), learned representations capture complex, distributed temporal and multisensor patterns more effectively than simple harmonics and summary statistics (Bengio et al., 2014).
+In particular, transformer-based and embedding-field models trained with self-supervision compress irregular, noisy multi-sensor time series into high-dimensional representations that are useful for downstream tasks (Tseng et al., 2024; Feng et al., 2025; Brown et al., 2025).
+Conceptually, the encoder in deep embedding models learns data-adaptive basis components and nonlinear combinations, whereas harmonic coefficients are fixed basis functions chosen a priori and may not adequately capture the subtle differences among very similar classes.
+Furthermore, self-supervised pre-training on large unlabeled data enhances embedding quality and improves label efficiency and generalization (Wang et al., 2022).
+These representational advantages likely explain the consistent gains in classification performance we observed across datasets of differing size and class balance.
+While deep learning has proven transformative in crop type mapping and land cover classification (Jakubik et al., 2023; Szwarcman et al., 2025; Wang et al., 2023), its application to tree species classification at NFI scale represents a relatively recent development.
+Our results demonstrate consistent improvements in evaluation metrics of 2–9 percentage points across datasets of varying class balance and complexity, suggesting that transformer-derived features capture phenological and structural variability more effectively than engineered metrics by Francini et al.
 (2024).
-Hence, we outline the methodological details in both complimentary approaches in the next two sections before detailing the experimental setup in section˜4.
-3.1 Harmonic and Seasonal Features
-We reproduced closely the classification approach of Francini et al.
-(2024), who recently published a new state-of-the-art classification approach for tree species mapping in the Netherlands.
-Their approach employ a RF Classifier (Breiman, 2001) with 500 trees that is fitted on 209 features containing Seasonal Medoid Statistics and Harmonic Features that we describe in the next two paragraphs.
-While Francini et al.
-(2024) used only features extracted from Sentinel-2 time series, we found experimentally (detailed later in section˜5.1 and table˜2) that including features from Sentinel-1 further slightly improves the results.
-Since we will also use Sentinel-1 to extract deep features (section˜3.2), we also include Sentinel-1 features to ensure a fair comparison.
-76 Seasonal Medoid Statistics
-In total, 12 bands from S1 and S2 are selected and additional 7 indices are calculated from S2 including NDVI, Normalized Burn Ratio (NBR), Enhanced Vegetation Index (EVI), and Tasseled Cap transformations: Brightness (TCB), Wetness (TCW), Greenness (TCG), Angle (TCA).
-This results in 19 bands, where for each band the median band value in each of the 4 seasons, winter (January, February, December), spring (March-May), summer (June-August), and autumn (September-November), the medoids values are extracted.
-These features describe the data through season-wise statistics, but do not capture the dynamic change in seasonality itself.
-133 Harmonic Features
-To capture seasonal changes, sine and cosine harmonics
-$P_{t}=\beta_{0}+\beta_{1}t+\beta_{2}\cos(2\pi\omega t)+\beta_{3}\sin(2\pi\omega t)$ (1)
-are fitted to the underlying monthly signal of each band according.
-The seven harmonic parameters $\beta_{0}$, $\beta_{1}$, $\beta_{2}$, $\beta_{3}$, $A$, $\phi$ for each of the 19 band are used as classification features alongside the residual RMSE: $\beta_{0}$ is the constant, $\beta_{1}$ is the time coefficient, and $\beta_{2}$ and $\beta_{3}$ are the frequency sine and cosine coefficients, respectively.
-To fit these four coefficients and to select $P_{t}$, the pixel $p$ harmonic values at time $t$, we used a least squares regression to fit Eq.
-1.
-$A=\sqrt{\beta_{2}^{2}+\beta_{3}^{2}}$ is the amplitude of the harmonic curve on the y-axis, $\phi=\arctan(\beta_{3}/\beta_{2})$ is the phase of the curve on the x-axis to the origin and $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(P_{t}-X_{t})^{2}}$ is root mean square error between $P_{t}$ and the actual pixel values $X_{t}$.
-The frequency $\omega$ = 1 (in years) and indicates one cycle per unit of time, initial guess values before fitting four coefficients were set as $\beta_{0}=0.1$, $\beta_{1}=0.1$, $\beta_{2}=0.4$, $\beta_{3}=0.4$.
-This approach was proposed by Francini et al. (2024) based on Sentinel-2 features. We added of Sentinel-1 features, as we found them to be slightly beneficial in accuracy (see section˜5.1).
-3.2 Deep Pre-trained Time Series Features
-Overview of Presto deep features.
-In contrast to harmonic features that fit mathematical curves to individual bands, we evaluate the effectiveness of deep features obtained by the pre-trained time series foundation model, which is Lightweight, Pre-trained Transformers for Remote Sensing Timeseries (Presto) developed and trained by Tseng et al.
+This performance gain supports the ongoing shift in remote sensing from task-specific feature engineering toward transferable, learned representations that can adapt to diverse ecological contexts without domain-specific tuning.
+6.3 Frozen vs fine-tune embeddings
+We include three deep embedding representations: AEF embeddings, TESSERA embeddings, and Presto embeddings (both frozen and fine‑tuned).
+These representations differ substantially in model scale, training data, and training methodology, which influences their performance in downstream species classification tasks.
+Across all datasets and evaluation metrics with the RF classifier, we find that embeddings extracted from the fine‑tuned Presto consistently improve downstream classification performance compared to using frozen Presto, AEF, or TESSERA embeddings with RF.
+Although AEF embeddings are trained with encoders ranging from 480M to 1B parameters on approximately 3B image observations, and TESSERA embeddings are trained with a 40M-parameter encoder on 0.8B time‑series pixels, both optimized to capture complex spatio‑temporal patterns across diverse land covers, these embeddings are produced from frozen (fixed) encoders.
+AEF’s model and weights are not publicly available.
+Therefore, its encoder cannot be fine-tuned for the specific forest species classification task evaluated here.
+Furthermore, while TESSERA is open and powerful, its large number of parameters requires substantial computational resources for fine-tuning and embedding extraction, which would likely overfit the small-scale training dataset.
+By contrast, Presto is trained with a 0.4M parameter encoder on 21.5M time‑series pixels, and its light-weighted pretrained model is openly available for fine‑tuning.
+Frozen Presto embeddings are extracted from the pretrained encoder without updating any model weights during fine‑tuning.
+As a relatively small model, the performance of frozen Presto embeddings is similar to or worse than that of harmonic + medoid features.
+However, when the same pretrained Presto encoder is further fine‑tuned on the downstream forest species classification objective, the resulting embeddings exhibit improved discriminative power compared to the other embeddings.
+This result highlights the advantage of task-specific adaptation, even though the fine-tuning process necessitates additional steps and computational effort.
+While foundation embeddings like AEF and TESSERA offer rich, general representations that serve as a solid baseline for many tasks, since they are pre-computed and easily extracted from coordinates without requiring raw time-series data processing, fine-tuned embeddings tailored to the specific classification task capture discriminative information more effectively.
+6.4 Neural network architectures for learned features
+The generally modest yet consistent performance differences observed between MLP and RF classifiers when trained on deep embeddings may also reflect differences in the training protocol and hyperparameter choices, not only intrinsic classifier inductive biases.
+In our experiments the RF models were trained using 70% of the labeled data, whereas the fine-tuning process for Presto (which includes MLP training) used a smaller effective training fraction because we split that 70% into 80%/20% train/validation folds for early stopping (i.e., 70% $\times$ 80% = 56% used for optimizer updates, with the remaining 14% used to select the lowest-validation-loss checkpoint).
+Using an internal validation split and early stopping is a common strategy to prevent overfitting in neural models, but it reduces the amount of data used for weight updates and can therefore affect measured performance, especially on small datasets (Mahsereci et al., 2017).
+We also adopted hyperparameter settings from Mouret et al.
+(2025).
+while those settings are reasonable starting points, hyperparameter sensitivity is known to substantially affect deep learning performance and optimal settings can depend on dataset size, class aggregation, and geographical region.
+Systematic hyperparameter search, nested cross-validation, or adaptive optimization (e.g., Bayesian) can therefore be important to conclusively ascribe performance differences to classifier architecture rather than to training procedure (Raiaan et al., 2024).
+Taken together, the observed small gains of MLP over RF are likely the result of both (i) the MLP’s ability to exploit non-linear structure in high-dimensional embeddings and (ii) procedural factors (reduced effective training set for MLP due to the validation split and reliance on transferred hyperparameters).
+To strengthen these conclusions, we recommend follow-up experiments such as increasing the effective training fraction (e.g., by using cross-validation or a larger outer training split), running a targeted hyperparameter search for the MLP on each dataset, and reporting stability across different validation splits and seeds.
+6.5 Computational efficiency and operational scalability
+From an operational perspective, this approach offers significant advantages for national-scale implementations.
+The computational efficiency is evidenced by Presto’s lightweight architecture, which enables fine-tuning on a single GPU or CPU (Tseng et al., 2024), making it accessible to organizations with limited computational resources.
+The pre-training process itself demonstrates remarkable efficiency, processing 21.5 million pixel time-series in only 2 hours 12 minutes per epoch (43 hours 15 minutes for 20 epochs), indicating that similar approaches could be readily applied to other countries’ NFI systems.
+For The Netherlands specifically, with its 36.4 million 10×10m pixels, large-scale pixel-label mapping appears computationally feasible using this methodology.
+6.6 Data quality, class balance, and spatial autocorrelation considerations
+The results in Table 4 indicate the role of class balance and sampling design in determining classification performance.
+Large performance gaps are observed between the original Dutch NFI datasets (SIMB and COMB) and the augmented SIBA dataset, despite using identical feature representations and classifiers.
+The class-balanced SIBA dataset achieves substantially higher performance (93.54 % F1 macro with the MLP classifier) compared to the imbalanced SIMB and COMB datasets (62.29 % and 50.76 % F1 macro, respectively).
+While SIBA also benefits from a larger sample size (13,790 samples versus approximately 1,460 for SIMB and COMB), the magnitude of improvement indicates that class balance is a primary driver of model performance rather than data volume alone.
+This is particularly evident for the macro-averaged F1 score, which is sensitive to underrepresented classes and more representative of operational performance across species.
+These findings suggest that future NFI data collection strategies could benefit from explicitly targeting underrepresented tree species to improve class balance.
+The data augmentation strategy proposed by Francini et al.
+(2024), which supplements field-based NFI observations with visually interpreted samples from high-resolution satellite imagery, provides a practical pathway to achieve this goal.
+From a spatial perspective, both the NFI and SIBA datasets incorporate measures to reduce spatial autocorrelation.
+While the original Dutch NFI sampling design selects only one plot per square kilometer, the SIBA dataset is augmented by delineating polygons around these NFI plots.
+Although SIBA enforces a minimum inter-sample distance of 15 m (Francini et al., 2024), this may not be sufficient to fully eliminate spatial autocorrelation in heterogeneous forest landscapes.
+In such environments, species composition and canopy structure can remain correlated over larger distances; for instance, Hermosilla et al.
+(2022) recommended a minimum sampling distance of 45 m.
+In spatial machine learning, failure to account for spatial dependence between training and testing samples can lead to overly optimistic performance estimates (Karasiak et al., 2022).
+Consequently, the use of a spatially independent validation dataset would be highly beneficial for ensuring a robust evaluation of NFI-scale applications.
+6.7 Ecological Interpretation of Classification Patterns
+Beyond quantitative performance metrics, the observed classification patterns have meaningful ecological implications that reflect the underlying biophysical and phenological characteristics of the plot types.
+While deep embeddings capture key phenological signals, they have intrinsic limits for discriminating species that share very similar seasonal trajectories: when two taxa leaf-out, peak and senesce at similar times and have comparable canopy optical properties, spectral-temporal separability is inherently low (Persson et al., 2018).
+This limitation can be mitigated by integrating complementary sensors that encode different trait axes such as canopy height, vertical structure from airborne LiDAR or GEDI, and spectral information (Lang et al., 2023; Wagner et al., 2025).
+Second, the superior performance of deep embeddings in our experiments is consistent with recent findings in the time-series remote sensing literature: self-supervised, pre-trained encoders learn high-dimensional phenological and contextual representations that generalize better across classes and recording conditions than a fixed set of harmonic + medoid statistics.
+These learned representations can therefore resolve subtle temporal patterns and interactions across sensors that hand-crafted features miss, explaining the empirical gains we report.
+Taken together, these interpretations not only provide deeper ecological context for the classification results, but also highlight how deep embeddings capture phenological patterns better than hand-crafted harmonic + medoid features.
+6.8 Limitations and future directions
+Several limitations merit consideration for operational deployment.
+First, deep embeddings were extracted from a model pre-trained on a limited temporal and spatial range (2020-2021 global data), and further domain adaptation might improve performance for underrepresented classes to country’s forests.
+Second, while MLPs demonstrated slightly superior performance in our controlled experiments, their generalization in operational settings with smaller, more imbalanced training sets or under different geographic and ecological conditions remains an open question.
+Future work should investigate robust training strategies, such as cross-validation, targeted hyperparameter optimization, and data augmentation, to better understand how neural classifiers trained on deep embeddings perform when labeled data are scarce or highly skewed.
+Such studies will help establish practical guidelines for deploying learned representations and downstream classifiers in real-world national forest inventory and monitoring applications.
+Third, our datasets derive exclusively from Dutch NFI data, which, while sufficiently heterogeneous within the context of temperate European forests, may not fully represent other bioclimatic zones.
+Applying deep embeddings to other datasets (Freudenberg et al., 2025a), regions, or ecosystems is essential for assessing robust generalization in NFI tree species classification.
+Future research directions include: (i) adaptation to new regions or tree species not present in the training data, testing the transferability of learned representations across different forest ecosystems; (ii) integration of additional modalities such as hyperspectral or LiDAR data where available, potentially capturing structural metrics that complement spectral-temporal patterns; and (iii) development of computational efficient uncertainty quantification methods to assess classification confidence for operational decision-making, (iv) inclusion of canopy height products (spaceborne GEDI or regional airborne LiDAR / CHMs) or/and higher spectral resolution data (hyperspectral or narrow-band indices) for the species that showed highest confusion with similar phenological patterns (e.g., Alnus, Fraxinus vs.
+Quercus).
+7 Summary and conclusions
+This study examined the efficacy of deep embeddings for tree species classification in National Forest Inventories, yielding several significant findings.
+First, deep embeddings extracted from pre-trained models consistently outperformed traditional methods employing harmonic and medoid predictors across varying dataset sizes and classification complexities.
+In particular, fine-tuned Presto embeddings achieve the best performance despite their small model size.
+This performance advantage was particularly evident with larger and more balanced training datasets, as demonstrated by the results from the augumented NFI dataset (SIBA), indicating superior scalability of deep learning approaches with increased high-quality data availability.
+Second, the MLP classifier demonstrated superior performance compared to RF in our experimental setting, corroborating previous findings (Mouret et al., 2025).
+Collectively, these results demonstrate the potential for implementing computationally efficient, large-scale tree species mapping within NFI systems.
+This research significantly advances our understanding of deep learning applications in forestry by having demonstrated the effectiveness of fine-tuning pre-trained deep embeddings for tree species classification, establishing a computationally efficient approach for improving NFI accuracy.
+This insight lays the groundwork for future developments in automated forest inventory systems and broader applications in environmental monitoring.
+CRediT authorship contribution statement
+Takayuki Ishikawa: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis, Conceptualization.
+Carmelo Bonannella: Writing – review & editing, Data curation, Supervision, Conceptualization.
+Bas J.
+W.
+Lerink Writing – review & editing, Data curation.
+Marc Rußwurm Writing – review & editing, Methodology, Supervision, Conceptualization.
+Declaration of competing interest
+C.B.
+reports employment with OpenGeoHub Foundation.
+Other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgement
+This research is based on data from the Dutch National Forest Inventory (NFI), commissioned by the Ministry of Agriculture, Nature and Food Quality of the Netherlands.
+We thank the Ministry and the agencies responsible for collecting and providing these nationally representative forest data, which underpin our analysis.
+C.B.
+(Carmelo Bonannella) was supported by the Open‑Earth‑Monitor Cyberinfrastructure project, funded by the European Union’s Horizon Europe research and innovation programme [grant agreement No.101059548].
+Data availability
+The source code is available on:
+Appendix A Appendix
+A.1 Implementation Details on Presto Embeddings
+Calculation of Presto deep embeddings.
+In contrast to harmonic features that fit mathematical curves to individual bands, we evaluate the effectiveness of deep embeddings obtained by the pre-trained time series foundation model, which is Lightweight, Pre-trained Transformers for Remote Sensing Timeseries (Presto) developed and trained by Tseng et al.
 (2024) for general-purpose remote sensing time series classification and regression.
 This model is based on transformer architecture and was originally pretrained on 21.5 million pixel time-series with 12-month contiguous intervals and static topographic and coordinates data.
-Each month’s composite satellite data included S1, S2 and its NDVI, ERA5, Dynamic World (DW), extracted between 2020-01-01 and 2021-12-31.
-The static SRTM and coordinates data are added to the time-series data as input of the Presto encoder.
+Each month’s composite satellite data included Sentinel-1 (S1), S-2 and its NDVI, European Centre for Medium-Range Weather Forecasts Re Analysis v5 (ERA5), Dynamic World (DW), extracted between 2020-01-01 and 2021-12-31.
+The static Shuttle Radar Topography Mission (SRTM) and coordinates data are added to the time-series data as input of the Presto encoder.
 The model was originally pretrained through Masked Auto Encoding (MAE) (He et al., 2022) reconstruct masked sequences across different satellite modalities.
-The advantages of this model include the ability to handle freely available multi-source and multi-temporal data even when some sources (e.g., ERA5) or temporal data (e.g., S2 values in November) are missing, and computational efficiency when processing large areas.
-We obtained 128 dimensional deep features from the encoder output.
+The advantages of this model include the ability to handle freely available multi-source and multi-temporal data even when some sources (e.g., ERA5) or temporal data (e.g., S-2 values in November) are missing, and computational efficiency when processing large areas.
+We obtained 128 dimensional deep embeddings from the encoder output.
 For this study, we used the pre-trained weights to initialize the Presto Encoder.
-Multi Layer Perceptron (MLP) classifier was attached for fine-tuning of Presto encoder to the downstream task of tree species classification, which is described in section˜4.
-We used the fine-tuned encoder to extract deep features from the test set for the comparison between Harmonic and Seasonal Features and Presto deep features.
-As shown in fig.˜2, we directly input 12-monthly time series ($T=12$ timesteps) for (i) 2 Sentinel-1 bands (S1), (ii) 10 Sentinel-2 bands (S2) and (iii) NDVI (NDVI), (iv) 2 ERA5 features (ERA5), (v) Dynamic world tensor with all tree classes value (equal to value of 1.
-DW).
+MLP classifier was attached for fine-tuning of Presto encoder to the downstream task of tree species classification, which is described in section˜4.
+We used the fine-tuned encoder to extract deep embeddings from the test set for the comparison between Harmonic and Seasonal Features and Presto deep embeddings.
+We directly input 12-monthly time series ($T=12$ timesteps) for (i) 2 Sentinel-1 bands (S-1), (ii) 10 Sentinel-2 bands (S-2) and (iii) NDVI, (iv) 2 ERA5 features (ERA5), (v) Dynamic World land cover class probabilities (DW).
 To every pixel-time series we appended two static-in-time products: (i) elevation and slope from the SRTM digital elevation model (Farr et al., 2007) (TG) and (ii) location coordinates (Loc) of each pixel as input tokens to fine-tuned Presto encoder.
-Location coordinates (Loc) is 3D static in time Cartesian coordinates computed from the latitude and longitude of the pixel’s geographical location.
+Location coordinates (Loc) are 3D static in time Cartesian coordinates computed from the latitude and longitude of the pixel’s geographical location.
 One training sample $\mathbf{x}$, comprising a pixel-timeseries $\mathbf{t}\in\mathbb{R}^{T\times 15}$, land cover classes $\mathbf{v}\in\mathbb{V}^{T\times 1}$, and static variables $\mathbf{s}\in\mathbb{R}^{1\times 5}$, is summarized as follows:
@@
   • 1.
-    $\mathbf{e}_{\text{S1}}$: S1 VV and VH bands.
+    $\mathbf{e}_{\text{S1}}$: S-1 VV and VH bands.
   • 2.
-    $\mathbf{e}_{\text{S2-RGB}}$: S2 RGB(B2, B3, B4) bands.
+    $\mathbf{e}_{\text{S2-RGB}}$: S-2 RGB(B2, B3, B4) bands.
   • 3.
-    $\mathbf{e}_{\text{S2-RE}}$: S2 Red Edge (B5, B6, B7) bands.
+    $\mathbf{e}_{\text{S2-RE}}$: S-2 Red Edge (B5, B6, B7) bands.
   • 4.
-    $\mathbf{e}_{\text{S2-NIR10}}$: S2 NIR (B8) 10m band.
+    $\mathbf{e}_{\text{S2-NIR10}}$: S-2 NIR (B8) 10m band.
   • 5.
-    $\mathbf{e}_{\text{S2-NIR20}}$: S2 NIR (B8A) 20m band.
+    $\mathbf{e}_{\text{S2-NIR20}}$: S-2 NIR (B8A) 20m band.
   • 6.
-    $\mathbf{e}_{\text{S2-SWIR}}$: S2 SWIR(B11, B12) bands.
+    $\mathbf{e}_{\text{S2-SWIR}}$: S-2 SWIR(B11, B12) bands.
   • 7.
-    $\mathbf{e}_{\text{NDVI}}$: Computed from S2 B4 and B8.
+    $\mathbf{e}_{\text{NDVI}}$: Computed from S-2 B4 and B8.
   • 8.
@@
   • 1.
-    Channel Group Encoding: A learnable encoding was added for each channel group $c\in\mathcal{C}=\{\text{S1},\text{S2 RGB},\ldots,\text{ERA5},\text{TG},\text{Loc}\}$, to capture source-specific information.
+    Channel Group Encoding: A learnable encoding was added for each channel group $c\in\mathcal{C}=\{\text{S-1},\text{S-2 RGB},\ldots,\text{ERA5},\text{TG},\text{Loc}\}$, to capture source-specific information.
 The final transformer input matrix $\mathbf{E}\in\mathbb{R}^{(T\cdot|\mathcal{C}_{\text{dynamic}}|+|\mathcal{C}_{\text{static}}|)\times d_{e}}$ is composed of:
@@
     $\mathbf{e}_{\text{Loc}}=h_{\text{Loc}}(s_{\text{Loc}})$
-Figure 2: Presto encoder feature extraction for downstream tasks. Figure modified from (Tseng et al., 2024).
-Data Normalization (for deep features).
+Data Normalization (for deep embeddings).
 The data normalization follows the original Presto setup (Tseng et al., 2024), with adjustments based on the observed value ranges for each source.
-Sentinel-1 VH and VV bands, originally ranging from –31 to 17 dB, are normalized by shifting +25 and dividing by 25, resulting in a range of –0.24 to 1.68.
+S-1 VH and VV bands, originally ranging from –31 to 17 dB, are normalized by shifting +25 and dividing by 25, resulting in a range of –0.24 to 1.68.
 Sentinel-2 spectral bands, with raw values from 10 to 15,769, are scaled by dividing by 10,000, yielding normalized values between 0.00 and 1.58.
@@
 SRTM elevation data, spanning –27 to 331 m, are divided by 2,000 to give a range of –0.01 to 0.17, while slope values from 0.0 to 39.3° are scaled by dividing by 50, yielding normalized values between 0.00 and 0.79.
-4 Experimental Setup
-This study compares classic machine learning with hand-crafted feature engineering against deep learning approaches for tree species classification.
-Our experimental design follows a systematic progression to evaluate different methodological approaches and their combinations.
-4.1 Comparative evaluation framework
-Our experimental design implements four distinct comparison scenarios to systematically evaluate feature representations and classifier architectures:
-  1. 1.
-    Sensor contribution analysis: S1 vs S2 vs S1+S2 using harmonic features with RF
-  2. 2.
-    Feature representation comparison: Harmonic features vs deep features using RF
-  3. 3.
-    Classifier architecture evaluation: RF vs MLP using deep features
-This systematic approach allows us to isolate the contributions of different methodological components while building toward the optimal configuration for operational NFI applications, extending the work of Francini et al.
-(2024) with comprehensive deep learning comparisons.
-4.2 Model fine-tuning
-The fine-tuning approach leverages the pre-trained Presto encoder’s learned representations while adapting the model to the specific characteristics of Dutch forest species.
-This transfer learning strategy balances computational efficiency with task-specific optimization.
-For fine-tuning the pre-trained models, the encoder block from the pre-trained model was reused and trained with newly attached Multi Layer Perceptron (MLP) classifier for the downstream task.
-This fine-tuning model is trained on the 1,479 and 13,790 samples for 7 aggregated groups classification task, and the 1,462 samples for the 13 dominant species classification task.
-Pre-trained encoder was set as trainable in this process.
-Based on the current state-of-the-art classifier of MLP on tree species classification (Mouret et al., 2024), this classifier architecture consists of a 3-layer MLP with no dropout (rate: 0.0).
-The layer configuration comprises 1024 nodes in the first layer, 512 nodes in the second layer, and 256 nodes for the last layer, with batch normalization and ReLU activation applied throughout.
-The training parameters are configured as follows: a learning rate of 0.0001 is used with a weight decay of 0.00746.
-The model is trained for a maximum of 100 epochs using a batch size of 64.
-The training process employs cross-entropy loss for optimization.
-The best fine-tuned models and its accuracy were selected based on best validation loss.
-4.3 Data splitting strategy
-Train-test split ratio is a critical factor in model performance on downstream tasks.
-If dataset size is large, 80/20 or higher ratios are commonly used (Rácz et al., 2021; Joseph, 2022).
-Based on our imbalanced dataset and previous research (Blickensdörfer et al., 2024), a 70/30 (approximately 2:1 ratio) train-test split was applied for fine-tuning and RF classification to increase model performance and reduce standard deviation on tree species classification over large areas for NFI.
-During fine-tuning and subsequent RF classification for features extracted from the fine-tuned model, the same training and test data were used.
-This practice ensures that extracted features are not trained on test data, which would cause data leakage.
-By adopting a class-per-split strategy, each class is equally split into train and test data at this ratio.
-In remote sensing machine learning, ensuring that the train-test split strategy does not introduce spatial autocorrelation is crucial for obtaining unbiased model performance estimates Karasiak et al.
-(2022).
-Our NFI data accommodates a sampling strategy which selects each plot with a density of 1 point per 100ha (Schelhaas et al., 2014).
-This means that each random point is designated per square kilometer.
-Additionally, the Francini data removed adjacent pixels with a minimum distance of 15m (Hermosilla et al., 2022; Francini et al., 2024).
-These sampling strategies ensure that the train-test split does not introduce strong spatial autocorrelation in the data; therefore, no additional train-test split strategy was applied in this study.
-4.4 Performance evaluation
-Model performance was evaluated using a comprehensive set of established accuracy metrics: per-class precision and recall, F1 scores, confusion matrices, and overall accuracy.
-While overall accuracy provides a general assessment of classification performance, it exhibits reduced sensitivity to class imbalance issues that are prevalent in ecological datasets.
-Therefore, the F1 score, which represents the harmonic mean of precision and recall, was employed as a supplementary metric to provide a more balanced evaluation of model performance when addressing the inherent class imbalances in tree species distribution data.
-Confusion matrices were used to visualize the model’s performance at the class level.
-These metrics are widely used in tree species classification studies (Goutte and Gaussier, ; Hermosilla et al., 2022; Francini et al., 2024; Blickensdörfer et al., 2024) and provide comprehensive assessment of classification performance across all species classes.
-4.5 Statistical validation
-To evaluate variability of model performance, each model was evaluated five times with different random seeds, which affect to train-test split and model initialization.
-The mean and standard deviation of the accuracy metrics were calculated for each model.
-This approach provides a more robust evaluation of model performance.
-5 Results
-5.1 Influence of multi-source input features
-Following the approach of Francini et al.
-(2024), we first evaluated the contribution of Sentinel-1 radar data to optical Sentinel-2 classification.
-Using RF with harmonic features, the addition of Sentinel-1 provided modest improvements: 0.4% for the 13-class COMB dataset and 1.4% for the 7-class SIMB dataset.
-These results justify the inclusion of Sentinel-1 features in our subsequent comparisons.
-We first evaluated whether combining optical and radar satellite data improves tree species classification using hand-crafted features in a RF classifier.
-Feature sets were derived from seasonal medoid composites and harmonic fits, using either Sentinel-2 (S2) alone or in combination with Sentinel-1 (S1).
-The goal was to assess whether radar backscatter, which captures structural and moisture-related canopy properties, offers complementary information to optical reflectance data.
-As shown in Table 2 the inclusion of S1 provided modest but consistent improvements across both datasets.
-For the 13-class dataset (COMB), accuracy increased from 66.34% to 66.71% when adding S1 to S2-derived features.
-In the 7-class dataset (SIMB), accuracy improved from 72.87% to 74.22%.
-These results suggest that the radar signal contributes marginally to distinguishing between species, especially under a more simplified class grouping.
-While the absolute gains remain small, incorporating S1 ensures a more robust feature representation, particularly in settings where optical signals may be limited by cloud cover or seasonal variability.
-The modest accuracy gains observed here are also useful to contextualize the more substantial differences in performance reported in the next sections, which stem from the feature representations rather than the data sources themselves.
-Table 2: Classification accuracy comparison across different feature combinations and satellite data sources in NFI data (in %). Bold values indicate best performance for each label type.
-Label Features Sentinel-1 Sentinel-2 S1-S2 Combined
-Band Acc. Band Acc. Band Acc.
-COMB (13 classes) Seasonal (S) S1 51.03 S2 64.03 S1-S2 64.52
-Harmonic (H) S1 55.04 S2 63.79 S1-S2 64.64
-All (S+H) S1 57.72 S2 66.34 S1-S2 66.71
-SIMB (7 classes) Seasonal (S) S1 55.60 S2 70.18 S1-S2 71.93
-Harmonic (H) S1 60.19 S2 71.39 S1-S2 71.52
-All (S+H) S1 62.75 S2 72.87 S1-S2 74.22
-5.2 Deep vs Harmonic features
-Table 3 demonstrates the superior performance of deep features over hand-crafted features when both are evaluated using RF classification.
-Deep features consistently outperform traditional harmonic and seasonal features across all three datasets, with F1 score improvements of 6-8 percentage points for the imbalanced datasets (COMB and SIMB) and over 10 percentage points for the balanced dataset (SIBA).
-Table 3: Classification performance comparison (in %) across feature types (deep vs.
-harmonic+medoid) and classifiers (MLP vs.
-RF).
+A.2 Details of methodology
+A.2.1 Data Splitting and training pipeline
+The detailed breakdown of class distribution across train, validation, and test splits is provided in table˜A.1.
+For fine-tuning Presto, we used MLP-Train data as training data for the MLP classifier and evaluated loss values based on MLP-validation data.
+Table A.1: Species/Group class distribution across train, validation, and test splits for three datasets.
+Dataset Class Train Test Total
+(train total) MLP-Train MLP-Val
+NFI (dominantSpecies) Alnus spp 21 16 5 9 30
+Betula spp 40 32 8 18 58
+Fagus spp 40 32 8 18 58
+Fraxinus spp 28 22 6 12 40
+Larix spp 39 31 8 17 56
+Other Pinus 62 49 13 27 89
+Other Quercus 23 18 5 10 33
+Other broadleaved 71 56 15 31 102
+Picea spp 46 36 10 20 66
+Pinus sylvestris 359 287 72 154 513
+Populus spp 50 40 10 22 72
+Pseudotsuga menziesii 62 49 13 28 90
+Quercus robur petraea 178 142 36 77 255
+Total 1,019 810 209 443 1,462
+NFI (Group_F) Beech 40 32 8 18 58
+Dark Conifer 112 89 23 48 160
+Larix 39 31 8 17 56
+Other Broadleaves 169 135 34 73 242
+Pinus 422 337 85 181 603
+Populus 50 40 10 22 72
+Quercus 201 160 41 87 288
+Total 1,033 824 209 446 1,479
+Francini (group2) Beech 1,379 1,103 276 591 1,970
+Dark Conifer 1,379 1,103 276 591 1,970
+Larix 1,379 1,103 276 591 1,970
+Other Broadleaf 1,379 1,103 276 591 1,970
+Pinus 1,379 1,103 276 591 1,970
+Populus 1,379 1,103 276 591 1,970
+Quercus 1,379 1,103 276 591 1,970
+Total 9,653 7,721 1,932 4,137 13,790
+A.3 Additional results
+A.3.1 Results of models using only S-1 and S-2 data for fine-tuning and embeddings extraction
+Our primary models incorporated all available satellite data as input for fine-tuning and extraction of deep embeddings from Presto.
+In this section, we present results from models using only S-1 and S-2 data as input for fine-tuning and embeddings extraction, excluding SRTM and ERA5 data.
+deep embeddings model performance
+TableA.2 reports classification performance in RF classifier when fine-tuning Presto using only S-1 and S-2 data for fine-tune and embeddings extraction, excluding auxiliary topographic (SRTM) and climatic (ERA5) inputs.
+Across all datasets, fine-tuned Presto embeddings substantially outperform harmonic + medoid features for all evaluation metrics, highlighting the strong representational capacity of deep embeddings even when limited to optical and radar data.
+These results indicate that S-1 + S-2 only fine-tuning Presto embeddings remains highly effective and robust.
+Table A.2: Classification performance comparison between fine-tuned Presto embeddings using only S-1 and S-2 inputs and harmonic + medoid features (in %).
+Results are reported for a RF classifier.
 Each model was run five times and the mean and standard deviation (±) are reported.
-Bold values indicate the best performance among MLP and RF per dataset type.
-Classification Results
-Dataset type Data size MLP Classifier
-(Deep features)
-RF Classifier
-(Deep features)
-RF Classifier
-(Harm.+med. Francini et al. (2024))
-Overall Accuracy SIMB (7 classes) 1479 77.13 ± 1.17 74.62 ± 1.93 70.58 ± 1.78
-COMB (13 classes) 1462 67.99 ± 1.33 66.77 ± 1.51 62.71 ± 1.51
-SIBA (7 classes) 13,790 98.21 ± 0.20 95.28 ± 0.60 84.31 ± 0.99
-F1 Score SIMB (7 classes) 1479 68.24 ± 2.85 60.54 ± 3.08 51.57 ± 3.36
-COMB (13 classes) 1462 52.45 ± 2.25 46.84 ± 2.65 38.54 ± 2.14
-SIBA (7 classes) 13,790 98.20 ± 0.20 95.27 ± 0.60 84.26 ± 0.99
-For direct comparison using identical S1+S2 sensor inputs (detailed in Appendix Table A.1), deep features show F1 score improvements of 12.3 percentage points (SIMB), 9.7 percentage points (COMB), and 10.7 percentage points (SIBA), confirming consistent performance advantages across different input configurations.
-We compared classification performance between traditional hand-crafted features and deep features extracted from a pre-trained transformer model.
-In both cases, a RF classifier was used to isolate the effect of the feature representation.
-As shown in Table 3, deep features extracted from the pre-trained Presto model consistently outperformed the harmonic and medoid predictors across all dataset.
-On the 7-class dataset (SIMB), models trained with deep features achieved an overall accuracy of 74.62% (±1.93), compared to 70.58% (±1.78) for the harmonic+medoid approach – an improvement of approximately 4 percentage points.
-The difference was even more pronounced when examining F1 scores, where deep features yielded 60.54% (±3.08) versus 51.57% (±3.36) for traditional features, representing an improvement of nearly 9 percentage points.
-In the more complex 13-classes case (COMB), deep features kept the advantage over traditional hand-crafted features, achieving an overall accuracy of 66.77% (±1.51) compared to 62.71% (±1.51).
-Similarly, the F1 score showed a substantial improvement from 38.54% (±2.14) with traditional features to 46.84% (±2.65) with deep features – a gain of more than 8 percentage points.
-This indicates that the deep learning approach handles the increased complexity of multi-class classification more effectively.
-The highest performance gains were observed in the Francini dataset (SIBA), where deep features achieved a remarkable overall accuracy of 95.28% (±0.60), representing an improvement of nearly 11 percentage points over the traditional features’ accuracy of 84.31% (±0.99).
-The F1 scores demonstrated comparable improvements, with deep features achieving 95.27% (±0.60) compared to 84.26% (±0.99) for traditional features.
-These results suggest that deep features extracted from the pre-trained transformer model are more effective at capturing species-specific spectral and temporal signatures than traditional hand-crafted features.
-The improvement is most pronounced in the SIBA dataset – where class separation is clearer due to the balanced classes and a higher sample size per class – but performance gains are evident across all dataset configurations examined in this study.
-To evaluate the influence of classifier architecture, we compared the Multi Layer Perceptron (MLP) classifier for tree species classification Mouret et al.
-(2024) with a RF, both trained on the same deep features extracted from the pre-trained Presto model.
-This isolates the effect of the classifier while keeping the input representation fixed.
-Across all datasets, the MLP classifier consistently outperformed the RF classifier in both overall accuracy and F1 score (Table 3).
-In the the 7-class dataset (SIMB), the MLP achieved an overall accuracy of 77.13% (±1.17) compared to 74.62% (±1.93) for the RF with a difference of approximately 2.5 percentage points.
-The performance gap was even more pronounced in terms of F1 score, with the MLP achieving 68.24% (±2.85) versus 60.54% (±3.08) for RF, representing a substantial improvement of nearly 8 percentage points.
-For the more challenging 13-class dataset (COMB), the MLP classifier maintained its advantage with an overall accuracy of 67.99% (±1.33) compared to 66.77% (±1.51) for the RF classifier.
-Similarly, the F1 score showed the MLP’s superior performance with 52.45% (±2.25) versus 46.84% (±2.65) for RF.
-The largest performance gain was observed in the balanced dataset (SIBA), where MLP reached 98.21% accuracy and 98.20% F1, compared to 95.28% and 95.27% with RF.
-Notably, the MLP classifier not only achieved higher mean performance values but also demonstrated more consistent results across repeated runs, as evidenced by the lower standard deviations in most cases.
-These results indicate that while both classifiers can effectively leverage deep features, the MLP classifier with pre-trained Presto offers superior performance for forest type classification tasks across datasets of varying complexity and size.
-The improved class separation suggests that MLPs better exploit species-specific spectral and temporal signatures embedded in the deep features—patterns that may be less accessible to tree-based models like RF, especially in more balanced or data-rich scenarios.
-The S1+S2 only comparison (detailed in Appendix Table A.2) confirms these findings with F1 score improvements of 6.7 percentage points for SIMB (70.57% vs 63.83%), 6.7 percentage points for COMB (54.91% vs 48.24%), and 3.1 percentage points for SIBA (98.07% vs 94.94%), demonstrating consistent MLP advantages across different input configurations.
-5.3 Class level performances
-To better understand class-level behavior, we visualized the confusion matrices of the best-performing model – an MLP fine-tuned on deep features – for all three datasets (COMB, SIMB, SIBA).
-These matrices are shown in Figure 3, with each cell reporting the number of predicted samples and the corresponding precision (i.e., percentage of correct predictions among all predicted samples for that class).
-In the 7-class NFI dataset (SIMB), the model performs strongly across most classes.
-Pinus shows high precision with 1,668 correctly predicted samples and a precision of 84.7%.
-Dark Conifer and Populus also achieve high precision with 1,448 samples at 73.5% and 1,833 samples at 93.0%, respectively.
-Broadleaved classes exhibit more overlap: Quercus (oak) has 1,478 samples with 75.0% precision, while Other Broadleaves and Fagus reach 71.6% and 91.6%, respectively.
-Confusion is mostly observed between Quercus, Populus, and Other Broadleaves, reflecting the spectral similarity within this group.
-In the more complex 13-class dataset (COMB), misclassifications are more pronounced, but key species maintain strong separation.
-Pinus sylvestris achieves 77.4% precision, Pseudotsuga menziesii reaches 83.3%, and Picea spp.
-also stands at 83.3%.
-Populus spp.
-improves to 73.9%, and Betula spp.
-achieves 100% precision, though on a smaller sample size.
-In contrast, Quercus robur/petraea reaches only 53.9%, and Other Broadleaves suffers from overlapping predictions with Quercus and Populus.
-Despite the increased class count, precision remains relatively high for most coniferous classes and for broadleaves with distinct phenology.
-In the balanced dataset (SIBA), the model achieves near-perfect class separation.
-All classes have 1,970 samples, and precision ranges between 84.0% (Larix) and 93.0% (Populus).
-Other classes include Fagus (91.6%), Pines (84.7%), Oak (75.0%), and Other Broadleaves (71.6%).
-The confusion is minimal and symmetrically distributed, confirming that the model benefits substantially from balanced data and larger sample size.
-Figure 3: Confusion matrices for the best performing models with deep features by MLP classifier.
-Simplified & Imbalanced (SIMB)(left), Complex & Imbalanced (COMB)(center), Simplified & Balanced (SIBA) (right)
-6 Discussion
-6.1 Multi-modal sensor integration
-Our findings on the modest but consistent benefits of combining Sentinel-1 radar with Sentinel-2 optical data align well with recent European tree species mapping studies.
-Blickensdörfer et al.
-(2024) demonstrated that Sentinel-1 SAR data effectively fills potential data gaps in cloud-affected regions, particularly important for national-scale applications where clear-sky observation density varies substantially.
-Similarly, Liu et al.
-(2023) confirmed that Sentinel-1 provides supplementary structural information to Sentinel-2 optical data, contributing to increased reliability of large-scale tree species mapping.
-Our results showing 0.4-1.4% accuracy improvements when adding Sentinel-1 to Sentinel-2 features corroborate these findings and provide confidence in the robustness of multi-modal approaches for operational NFI applications.
-6.2 Deep vs hand-crafted features: an emerging paradigm
-The transition from hand-crafted to learned feature representations marks a significant shift in remote sensing methodology.
-While deep learning has proven transformative in crop type mapping and land cover classification (Jakubik et al., 2023; Szwarcman et al., 2025; Wang et al., 2023), its application to tree species classification at NFI scale represents a relatively recent development.
-Our results demonstrate consistent improvements in F1 scores of 6-12 percentage points across datasets of varying class balance and complexity, suggesting that transformer-derived features capture phenological and structural variability more effectively than engineered metrics.
-This performance gain supports the ongoing shift in remote sensing from task-specific feature engineering toward transferable, learned representations that can adapt to diverse ecological contexts without domain-specific tuning.
-6.3 Neural network architectures for learned features
-The systematic superiority of MLP classifiers over RF when both were trained on deep features reflects broader patterns observed in high-dimensional, learned feature spaces.
-Neural architectures tend to better exploit the latent structure present in deep representations, where complex non-linear relationships between features may not be effectively captured by tree-based models.
-This finding aligns with recent advances in representation learning, where the pairing of learned features with compatible downstream architectures has proven critical for optimal performance (Mouret et al., 2024).
-6.4 Computational efficiency and operational scalability
-From an operational perspective, this approach offers significant advantages for national-scale implementations.
-The computational efficiency is evidenced by Presto’s lightweight architecture, which enables fine-tuning on a single GPU or CPU (Tseng et al., 2024), making it accessible to organizations with limited computational resources.
-The pre-training process itself demonstrates remarkable efficiency, processing 21.5 million pixel time-series in only 2 hours 12 minutes per epoch (43 hours 15 minutes for 20 epochs), indicating that similar approaches could be readily applied to other countries’ NFI systems.
-For The Netherlands specifically, with its 36.4 million 10×10m pixels, large-scale pixel-label mapping appears computationally feasible using this methodology.
-6.5 Data quality and class balance considerations
-Our comparative analysis across datasets of varying quality and balance reveals critical insights for operational NFI implementations.
-The stark performance differences between the NFI datasets (COMB and SIMB) and the Francini dataset (SIBA) underscore the fundamental importance of training data characteristics for deep learning approaches.
-The Francini dataset’s superior performance (98.2% F1 score) compared to the imbalanced NFI datasets (52.5-68.2% F1 scores) primarily reflects the impact of class balance rather than dataset size alone.
-While the Francini dataset benefits from both larger sample size (13,790 vs 1,462-1,479) and better class balance, the magnitude of improvement suggests that addressing class imbalance represents a critical optimization opportunity for operational systems.
-This finding has important implications for NFI data collection strategies.
-Resources may be more effectively allocated to targeted collection for underrepresented species classes.
-The approach demonstrated by Francini et al.
-(2024), which supplements ground-truth NFI data with additional samples through visual interpretation of high-resolution satellite imagery, offers a practical pathway for addressing these imbalances.
-Tools such as Collect Earth Online (Saah et al., 2019) provide accessible platforms for additional data collection.
-6.6 Limitations and future directions
-Several limitations merit consideration for operational deployment.
-First, deep features were extracted from a model pre-trained on a limited temporal and spatial range (2020-2021 global data), and further domain adaptation might improve performance for underrepresented classes to country’s forests.
-Second, while MLPs demonstrated superior performance in our controlled experiments, their generalization in operational settings with smaller or more imbalanced training sets requires further investigation.
-Future research directions include: (i) adaptation to new regions or tree species not present in the training data, testing the transferability of learned representations across different forest ecosystems; (ii) integration of additional modalities such as hyperspectral or LiDAR data where available, potentially capturing structural metrics that complement spectral-temporal patterns; and (iii) development of computational efficient uncertainty quantification methods to assess classification confidence for operational decision-making.
-7 Summary and conclusions
-This study examined the efficacy of deep learning approaches for tree species classification in National Forest Inventories, yielding several significant findings.
-First, deep features extracted from pre-trained Presto models consistently outperformed traditional methods employing harmonic and medoid predictors across varying dataset sizes and classification complexities.
-This performance advantage was particularly evident with larger and more balanced training datasets, as demonstrated by the results from the Francini dataset, indicating superior scalability of deep learning approaches with increased high-quality data availability.
-Second, the MLP classifier demonstrated superior performance compared to RF in our experimental setting, corroborating previous findings (Mouret et al., 2024).
-Collectively, these results demonstrate the potential for implementing computationally efficient, large-scale tree species mapping within NFI systems.
-This research significantly advances our understanding of deep learning applications in forestry by having demonstrated the effectiveness of fine-tuning pre-trained deep learning models for tree species classification, establishing a computationally efficient approach for improving NFI accuracy.
-This insight lays the groundwork for future developments in automated forest inventory systems and broader applications in environmental monitoring.
-CRediT authorship contribution statement
-Takayuki Ishikawa: Writing – review & editing, Writing – original draft, Visualization, Methodology, Investigation, Formal analysis, Conceptualization.
-Carmelo Bonannella: Writing – review & editing, Supervision, Conceptualization.
-Bas J.
-W.
-Lerink Writing – review & editing, Data curation.
-Marc Rußwurm Writing – review & editing, Supervision, Conceptualization.
-Declaration of competing interest
-C.B.
-reports employment with OpenGeoHub Foundation.
-Other authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Acknowledgement
-This research is based on data from the Dutch National Forest Inventory (NFI), commissioned by the Ministry of Agriculture, Nature and Food Quality of the Netherlands.
-We thank the Ministry and the agencies responsible for collecting and providing these nationally representative forest data, which underpin our analysis.
-C.B.
-(Carmelo Bonannella) was supported by the Open-Earth-Monitor Cyberinfrastructure project, funded by the European Union’s Horizon Europe research and innovation programme [grant agreement No.101059548].
-Data availability
-The source code is available on:
-Appendix A Appendix
-A.1 Additional results
-A.1.1 Results of models using only S1 and S2 data for fine-tuning
-Our primary models incorporated all available satellite data as input for both the second stage of pre-training and fine-tuning to extract deep features.
-In this section, we present results from models using only S1 and S2 data as input for fine-tuning and feature extraction, excluding SRTM and ERA5 data.
-Deep features model performance
-Table A.1 demonstrates that deep features extracted by pre-trained Presto from only S1 and S2 data outperformed harmonic and medoid features in the RF classifier.
-These results were comparable to models that utilized all satellite data as input.
-Table A.1: Classification performance comparison between deep features and harmonic+medoid features (in %).
+Bold values indicate the best performance between the models per dataset type.
+RF
+Dataset type Data size Fine-tune Presto
+embeddings
+Harm.+med.
+features
+((Francini et al., 2024))
+Overall Accuracy SIMB (7 classes) 1,479 74.01 ± 1.83 69.48 ± 0.73
+COMB (13 classes) 1,462 65.42 ± 1.92 62.80 ± 0.91
+SIBA (7 classes) 13,790 92.31 ± 0.18 84.47 ± 0.41
+F1 Macro SIMB (7 classes) 1,479 60.00 ± 2.12 51.03 ± 1.71
+COMB (13 classes) 1,462 44.67 ± 3.25 38.95 ± 2.02
+SIBA (7 classes) 13,790 91.95 ± 0.24 84.40 ± 0.40
+F1 Weighted SIMB (7 classes) 1,479 72.96 ± 1.87 66.98 ± 1.06
+COMB (13 classes) 1,462 61.32 ± 2.10 57.13 ± 1.25
+SIBA (7 classes) 13,790 92.23 ± 0.20 84.40 ± 0.40
+Table A.3 compares the performance of MLP and RF classifiers using fine-tuned Presto embeddings when only S-1 and S-2 data are provided as input.
+The MLP classifier consistently outperforms RF on the SIMB and COMB datasets across all metrics, with particularly notable improvements in F1 macro and weighted scores for the more class-imbalanced COMB dataset.
+In contrast, performance differences between the two classifiers are marginal for the SIBA dataset, where both models achieve similarly high accuracy and F1 scores.
+This suggests that classifier choice has a stronger impact under reduced input modalities and class imbalance, while its effect diminishes for larger and more balanced datasets.
+Table A.3: Classification performance comparison between MLP and RF classifiers using fine-tuned Presto embeddings with only S-1 and S-2 inputs (in %).
 Each model was run five times and the mean and standard deviation (±) are reported.
-Bold values indicate the best performance between deep features extracted from Presto and harmonic+medoid features per dataset type.
-RF
-Dataset type Data size Deep feat.
-((Tseng et al., 2024))
-Harm.+med.
-((Francini et al., 2024))
-Overall Accuracy NFI data (7 classes) 1480 75.47 ± 1.15 70.58 ± 1.78
-NFI data (13 classes) 1462 67.77 ± 2.38 62.71 ± 1.51
-Francini data (7 classes) 13,790 94.96 ± 0.58 84.31 ± 0.99
-F1 Score NFI data (7 classes) 1480 63.83 ± 1.21 51.57 ± 3.36
-NFI data (13 classes) 1462 48.24 ± 4.31 38.54 ± 2.14
-Francini data (7 classes) 13,790 94.94 ± 0.58 84.26 ± 0.99
-Table A.2 shows that the MLP classifier consistently achieved higher accuracy than the RF classifier, even when using only S1 and S2 data as input for the pre-trained Presto model.
-Table A.2: Classification performance comparison between MLP and RF classifiers with deep features (in %).
+Bold values indicate the best performance between the MLP and RF classifiers per dataset type.
+Fine-tune Presto embeddings
+Dataset type Data size MLP RF
+Overall Accuracy SIMB (7 classes) 1,479 74.98 ± 1.78 74.01 ± 1.83
+COMB (13 classes) 1,462 67.09 ± 0.60 65.42 ± 1.92
+SIBA (7 classes) 13,790 92.23 ± 0.48 92.31 ± 0.18
+F1 Macro SIMB (7 classes) 1,479 65.09 ± 2.97 60.00 ± 2.12
+COMB (13 classes) 1,462 50.82 ± 3.87 44.67 ± 3.25
+SIBA (7 classes) 13,790 91.91 ± 0.54 91.95 ± 0.24
+F1 Weighted SIMB (7 classes) 1,479 74.49 ± 1.88 72.96 ± 1.87
+COMB (13 classes) 1,462 65.00 ± 1.28 61.32 ± 2.10
+SIBA (7 classes) 13,790 92.20 ± 0.49 92.23 ± 0.20
+Comparison between all bands and only S-1 and S-2 bands
+Table A.4 compares fine-tuned Presto embeddings obtained using all available satellite bands with those using only S-1 and S-2 data, evaluated with the MLP classifier.
+Overall, performance differences between the two band configurations are small across datasets, indicating that excluding auxiliary climatic (ERA5) and topographic (SRTM) information does not substantially degrade classification performance for this task.
+For the SIMB dataset, using only S-1 and S-2 bands yields slightly higher accuracy and F1 scores, whereas for the COMB dataset both configurations perform comparably.
+In contrast, for the SIBA dataset, which is larger and more balanced, the use of all available bands results in marginally higher performance across all metrics.
+These results suggest that optical and radar data alone are sufficient to capture most discriminative information, while auxiliary bands provide limited additional benefit that appears to be dataset-dependent.
+Table A.4: Classification performance comparison between using all available bands and using only S-1 and S-2 bands for fine-tuned Presto embeddings with an MLP classifier (in %).
 Each model was run five times and the mean and standard deviation (±) are reported.
-Bold values indicate the best performance between MLP and RF classifier per dataset type.
-Results
-Dataset type Data MLP RF
-size classifier classifier
-Overall Accuracy NFI data (7 classes) 1480 78.79 ± 1.01 75.47 ± 1.15
-NFI data (13 classes) 1462 69.39 ± 1.65 67.77 ± 2.38
-Francini data (7 classes) 13,790 98.08 ± 0.09 94.96 ± 0.58
-F1 Score NFI data (7 classes) 1480 70.57 ± 2.20 63.83 ± 1.21
-NFI data (13 classes) 1462 54.91 ± 4.29 48.24 ± 4.31
-Francini data (7 classes) 13,790 98.07 ± 0.09 94.94 ± 0.58
-Comparison between all bands and only S1 and S2 bands
-Table A.3 compares models using all satellite data against those using only S1 and S2 data with MLP classifier.
-Interestingly, models with only S1 and S2 data achieved higher accuracy with NFI data, while no significant difference was observed with Francini data.
-Table A.3: Classification performance comparison between using all available bands versus using only S1 and S2 bands (in %).
-Each model was run five times with mean and standard deviation (±) reported.
-Bold values indicate best performance between the two band configurations for each dataset type.
-MLP classifier
-Dataset type Data size All bands S1S2 only
-Overall Accuracy NFI data (7 classes) 1480 77.13 ± 1.17 78.79 ± 1.01
-NFI data (13 classes) 1462 67.99 ± 1.33 69.39 ± 1.65
-Francini data (7 classes) 13,790 98.21 ± 0.20 98.08 ± 0.09
-F1 Score NFI data (7 classes) 1480 68.24 ± 2.85 70.57 ± 2.20
-NFI data (13 classes) 1462 52.45 ± 2.25 54.91 ± 4.29
-Francini data (7 classes) 13,790 98.20 ± 0.20 98.07 ± 0.09
+Bold values indicate the best performance between the two band configurations for each dataset type.
+MLP for fine-tune embeddings
+Dataset type Data size All bands S-1 + S-2 only
+Overall Accuracy SIMB (7 classes) 1,479 73.67 ± 1.29 74.98 ± 1.78
+COMB (13 classes) 1,462 67.40 ± 1.47 67.09 ± 0.60
+SIBA (7 classes) 13,790 93.54 ± 1.27 92.23 ± 0.48
+F1 Macro SIMB (7 classes) 1,479 62.29 ± 4.60 65.09 ± 2.97
+COMB (13 classes) 1,462 50.76 ± 2.23 50.82 ± 3.87
+SIBA (7 classes) 13,790 93.54 ± 1.26 91.91 ± 0.54
+F1 Weighted SIMB (7 classes) 1,479 73.05 ± 1.52 74.49 ± 1.88
+COMB (13 classes) 1,462 65.51 ± 1.29 65.00 ± 1.28
+SIBA (7 classes) 13,790 93.54 ± 1.26 92.20 ± 0.49
+Confusion Matrix Comparison
+Sets of confusion matrices for SIBA and COMB datasets show per-class accuracy improvements from deep embeddings, especially with fine-tuned Presto and the MLP classifier.
+Figure 3 presents confusion matrices for the COMB dataset across six model configurations, combining different feature representations (fine-tuned deep embeddings, frozen deep embeddings, harmonic + medoid feature) and classifiers (MLP and Random Forest).
+This appendix provides a comparative description of class-level behavior across these configurations, complementing the detailed discussion of the fine-tuned MLP results in Section 5.3.
+Across all models, broadleaved species exhibit substantially higher confusion than conifer species.
+In particular, Q.
+robur/petraea consistently acts as a dominant prediction class, absorbing misclassifications from several less frequent broadleaved species such as Alnus spp., Fraxinus spp., and Populus spp..
+This pattern is visible irrespective of classifier choice, indicating that class imbalance and overlapping spectral–phenological characteristics dominate model behavior more strongly than the specific learning algorithm.
+Comparing classifiers, RF models generally show more fragmented confusion patterns than the fine-tuned MLP.
+While RF-based models correctly identify dominant classes such as P.
+sylvestris and Q.
+robur/petraea, they exhibit reduced sensitivity to minority classes, with many rare species being almost entirely misclassified into majority classes.
+This effect is particularly evident for Alnus spp.
+and Fraxinus spp., which achieve very low recall across all RF configurations.
+Feature representation also plays a key role.
+Fine-tuned deep embeddings yield the most diagonally concentrated confusion matrix, especially for conifer species.
+P.
+sylvestris and P.
+menziesii are consistently well separated under fine-tuned representations, whereas frozen deep features (Presto, AEF, and TESSERA) show increased confusion between conifer taxa, particularly between Pinus spp.
+and P.
+sylvestris.
+This suggests that task-specific fine-tuning is critical for capturing subtle intra-genus differences.
+Models based on harmonic and medoid features display the weakest class separation overall.
+While they retain some ability to discriminate coarse functional groups (broadleaved vs.
+conifer), fine-grained discrimination among broadleaved species is limited.
+Finally, large-scale pretrained representations (AEF and TESSERA) exhibit mixed behavior.
+While they perform competitively for dominant classes, they do not consistently improve minority-class recognition in the absence of fine-tuning.
+In several cases, their confusion patterns resemble those of frozen Presto embeddings, indicating that representation scale alone does not resolve class imbalance or inter-species similarity issues in the COMB dataset.
+Overall, the confusion matrices for COMB reinforce three key observations: (i) class imbalance strongly shapes prediction outcomes across all models, (ii) fine-tuning deep embeddings improves intra-group species discrimination, and (iii) classifier choice has a secondary effect compared to data characteristics and feature representation.
+These findings support the conclusions drawn in the main text regarding the importance of data quality and class balance for operational NFI species mapping.
+Figure 3: Confusion matrices for Complex & Imbalanced (COMB) datasets.
+Top left: Fine- tune Presto with MLP Classifier, Top right: Presto (Frozen Presto features), Middle left: Fine-tune Presto (Presto features), Middle right: Harmonic + medoid feature, Bottom left: AEF, Bottom right: TESSERA.
+Each cell has average count and recall percentage.
+Figure 4 shows confusion matrices for the SIBA dataset across six model configurations, combining different feature representations and classifiers.
+In contrast to the COMB dataset, SIBA exhibits markedly stronger class separability across all models, reflecting its larger sample size and substantially improved class balance.
+Across configurations, most classes achieve high diagonal dominance, indicating strong recall and precision for both broadleaved and conifer species groups.
+Fine-tuned Presto with an MLP classifier yields near-optimal performance for all classes, with precision values exceeding 90 % for most species and limited off-diagonal confusion.
+This confirms that, under favorable data conditions, fine-tuned deep embeddings can effectively capture inter-species differences in seasonal and structural signals.
+RF classifiers show slightly reduced performance compared to the fine-tuned MLP but retain strong class-level behavior.
+In particular, dominant classes such as Beech, Larix, and Populus are consistently well identified across all feature representations.
+Residual confusion is primarily concentrated among broadleaved classes, especially between Other broadleaf and Quercus, suggesting remaining spectral and phenological overlap even in a balanced dataset.
+Frozen deep feature representations (Presto, AEF, and TESSERA) display broadly similar confusion patterns.
+While they maintain high accuracy for majority classes, they exhibit increased confusion for classes with overlapping ecological traits.
+For instance, Dark conifer shows increased misclassification into Pinus and Larix under frozen representations, indicating that fine-tuning remains important for resolving subtle intra-group distinctions.
+Models based on harmonic and medoid features show the weakest overall class separation, although their performance remains notably stronger than in the COMB dataset.
+This suggests that the improved balance and density of the SIBA dataset partially compensate for the lower representational capacity of hand-crafted temporal features.
+Nonetheless, confusion among broadleaved species remains more pronounced than for conifers, consistent with known challenges in deciduous species discrimination.
+Overall, the SIBA confusion matrices demonstrate that improved class balance and increased sample size substantially reduce misclassification across all modeling approaches.
+While fine-tuned deep embeddings provide the most consistent and robust class-level performance, the reduced confusion observed even for simpler models highlights the critical role of data quality in species-level forest classification.
+These results support the conclusion that targeted data augmentation and balancing strategies can be as influential as model architecture in improving operational NFI mapping performance.
+Figure 4: Confusion matrices for Simplified & Imbalanced (SIMB) datasets.
+Top left: Fine- tune Presto with MLP Classifier, Top right: Presto (Frozen Presto features), Middle left: Fine-tune Presto (Presto features), Middle right: Harmonic + medoid feature, Bottom left: AEF, Bottom right: TESSERA.
+Each cell has average count and recall percentage.